### v2

# Temporal Conformal Prediction (TCP): A Distribution-Free Statistical and Machine Learning Framework for Adaptive Risk Forecasting

Abstract
We propose Temporal Conformal Prediction (TCP), a principled framework for constructing well-calibrated prediction intervals for non-stationary time series. TCP integrates a machine learning-based quantile forecaster with an online conformal calibration layer. This layer’s thresholds are updated via a modified Robbins-Monro scheme, allowing the model to dynamically adapt to volatility clustering and regime shifts without rigid parametric assumptions. We benchmark TCP against GARCH, Historical Simulation, and static Quantile Regression across diverse financial assets. Our empirical results reveal a critical flaw in static methods: while sharp, Quantile Regression is poorly calibrated, systematically over-covering the nominal 95% target. In contrast, TCP’s adaptive mechanism actively works to achieve the correct coverage level, successfully navigating the coverage-sharpness tradeoff. Visualizations during the 2020 market crash confirm TCP’s superior adaptive response, and a comprehensive sensitivity analysis demonstrates the framework’s robustness to hyperparameter choices. Overall, TCP offers a practical and theoretically-grounded solution to the central challenge of calibrated uncertainty quantification for time series under distribution shift, advancing the interface between statistical inference and machine learning.
Keywords: Temporal Conformal Prediction, Quantile Regression, Machine Learning, Statistical Learning, Value-at-Risk, Financial Time Series, Risk Management.
MSC 2020 Subject Classification: 62G08, 62M10, 62P05, 91G70, 68T05
1.   Introduction
Financial risk estimation is more than a regulatory checkbox. It’s a foundational tool for ensuring market stability and investor confidence (Markowitz, 1952; J.P. Morgan, 1996). Yet, when markets enter turbulent regimes, traditional risk models often fall short. Value-at-Risk (VaR) and Expected Shortfall (ES) methods, for instance, typically assume Gaussian or other parametric return distributions (Hull & White, 1998), which systematically underestimate extreme losses. Events like the 2008 global financial crisis (Brunnermeier, 2009) and the COVID-19 crash (Baker et al., 2020) exposed the fragility of these assumptions, especially in the tails.
The core challenge is that financial returns violate the independent and identically distributed (i.i.d.) assumption that underpins many statistical learning techniques. Real-world returns are heteroskedastic, exhibit regime shifts, and often have heavy-tailed distributions (Cont, 2001). GARCH-type models (Engle, 1982; Bollerslev, 1986) can capture some of this volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
In recent years, conformal prediction has gained traction as a powerful distribution-free framework for quantifying uncertainty. Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk et al., 2005). However, these guarantees rest on the assumption of exchangeability, a condition often violated by time series data with temporal dependencies. Several recent works attempt to address this by extending conformal methods to sequential or temporally dependent settings (Xu & Xie, 2021; Stankeviciute et al., 2021), but a robust and practical solution for financial markets remains an open challenge.
This paper closes that gap. We introduce Temporal Conformal Prediction (TCP), a real-time, adaptive framework for constructing well-calibrated prediction intervals for financial time series. TCP combines a modern machine learning-based quantile forecaster with an online conformal calibration layer governed by a modified Robbins-Monro scheme. Crucially, this architecture bridges statistical inference and machine learning, combining the theoretical rigor of conformal methods with the flexibility of a data-driven approach that can adapt to non-stationarity, volatility clustering, and abrupt market shifts.
We benchmark TCP across three major asset classes, equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), against established models like GARCH, Historical Simulation, and standard Quantile Regression (QR). Our results highlight a critical flaw in static ML approaches: while QR produces sharp intervals, it is poorly calibrated and systematically over-covers the target 95% confidence level. TCP, by contrast, demonstrates superior adaptive capabilities, adjusting its intervals in response to market volatility to better align with the desired coverage rate. This adaptive property, which we visualize during the COVID-19 crash, makes TCP a more reliable and principled framework for real-world risk management. Furthermore, a comprehensive sensitivity analysis underscores the robustness of our framework to key hyperparameter choices.
We organize the paper as follows. Section 2 surveys related work. Section 3 establishes our mathematical notation. Section 4 lays out the TCP theory. Section 5 details our empirical setup, including the model architectures and evaluation framework. Section 6 presents our main findings, including the benchmark comparisons, visualizations, and hyperparameter sensitivity analysis. Finally, Section 7 wraps up with practical takeaways and future directions.
2.   Related Work
Before detailing our proposed TCP framework, we situate our work within the existing literature. This section reviews three key research areas that inform our approach: traditional financial risk models, the evolution of time series forecasting with machine learning, and the theoretical foundations of conformal prediction.
Traditional Financial Risk Models. Quantitative risk management traces its roots to Markowitz’s portfolio theory (Markowitz, 1952) and was formalized around the Value-at-Risk concept with J.P. Morgan’s RiskMetrics in 1996 (Morgan, 1996). In practice, three main VaR estimation paradigms have dominated: parametric approaches like GARCH-style models, which assume a specific return distribution; non-parametric Historical Simulation, which relies on past returns; and Monte Carlo methods, which simulate future scenarios. Beyond VaR, Conditional Value-at-Risk (CVaR) was proposed as a coherent alternative (Rockafellar & Uryasev, 2000). However, a common weakness is that these traditional approaches often fail backtests during market crises (Kupiec, 1995; Christoffersen, 1998), highlighting their blind spot when tail events dominate.
Advanced Time Series and Machine Learning Models. To better capture the dynamics of financial data, the GARCH family of models became foundational for modeling time-varying variance and volatility clustering (Engle, 1982; Bollerslev, 1986). Extreme Value Theory (EVT) offered a principled way to model the tails of distributions directly (Pickands, 1975; McNeil & Frey, 2000). More recently, machine learning has injected fresh momentum into risk estimation. Quantile regression (Koenker & Bassett, 1978) directly targets conditional quantiles without assuming a distribution, with modern variants including gradient-boosting forests (Meinshausen, 2006). Deep learning, especially LSTM networks, has also shown promise in capturing complex temporal dependencies (Fischer & Krauss, 2018).
Conformal Prediction Theory. Bridging these worlds is conformal prediction (Vovk et al., 2005), a framework that provides distribution-free, finite-sample prediction intervals. However, classical conformal methods assume exchangeability, a condition that breaks down in time series. Recent advances have sought to address this limitation by introducing adaptive calibration to handle distribution shift (Gibbs & Candès, 2021) and by developing extensions specifically for temporal data (Xu & Xie, 2021; Stankeviciute et al., 2021). Our work builds on these foundations, proposing a practical and robust adaptive conformal framework tailored specifically for the challenges of financial risk forecasting.
3.   Notation
Having reviewed the relevant literature, we now establish the mathematical notation that will be used throughout the remainder of the paper. A clear notational framework is essential for formally developing our proposed method in the subsequent sections.
We denote the price of a financial asset at time $t$ as $P_{t}$, and its corresponding daily log-return as $r_{t}$. The primary objective is to construct a $(1-\alpha)$ prediction interval, denoted by $[\ell_{t+1},u_{t+1}]$, for the next return $r_{t+1}$, where $\alpha\in(0,1)$ is the specified miscoverage rate. The feature space for our models includes lagged returns and realized volatility $\sigma_{t}$.
Learning and Prediction: For a general learning problem with $n$ observations, we consider feature-label pairs as $(X_{i},Y_{i})$ and a generic point prediction model as $\widehat{f}(\cdot)$. Our time series evaluation runs for $T$ total time steps. We use $\mathcal{F}_{t}$ to represent the filtration (information set) available up to time $t$. The true, unknown $\tau$-quantile of the return distribution is $q_{\tau}$. Our quantile regression model, $\mathcal{F}_{\tau}$, produces a data-driven estimate of this quantile, denoted $\widehat{q}_{\tau}$.
Conformal Prediction: Our conformal prediction framework relies on a non-conformity score function, $A(\cdot,\cdot)$, which produces a score $\alpha_{i}$ for each observation. For our method, we use the model residuals as scores, $\varepsilon_{i}$. This process generates a conformal prediction set, $\Gamma_{\alpha}$. The method’s local adaptation is managed using a rolling window of size $w$, which informs the adaptive conformal threshold at time $t$, denoted $C_{t}$.
Adaptive Calibration: The online adaptive calibration is driven by the coverage error at time $t$, defined as $e_{t}=\mathbb{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$. The threshold is updated using a learning rate $\gamma_{t}$, which is itself controlled by hyperparameters $\gamma_{0},\lambda,$ and $\beta$.
Benchmark Models: For the benchmark models, we denote the conditional variance from the GARCH model as $\hat{\sigma}_{t}^{2}$, which is governed by parameters $\omega,\alpha_{\text{GARCH}},$ and $\beta_{\text{GARCH}}$. For Historical Simulation, we use $\bar{r}$ to represent the mean return over a rolling window.
[…]
4.   Mathematical Background
This section lays the theoretical groundwork for our proposed method. We begin by formally defining the problem of prediction interval forecasting, then review the principles of classical conformal prediction and its limitations, and finally develop the core mathematical components of our Temporal Conformal Prediction (TCP) framework and its adaptive calibration mechanism.
4.1 Problem Formulation
We begin by framing the central task of this paper. We consider a univariate financial time series of daily log-returns $\{r_{1},r_{2},\dots,r_{t}\}$, where $r_{t}=\log(P_{t}/P_{t-1})$ and $P_{t}$ is the asset price at time $t$. Our objective is to construct a prediction interval $[\ell_{t+1},u_{t+1}]$ for the next return $r_{t+1}$ that satisfies the nominal coverage property $\Pr(r_{t+1}\in[\ell_{t+1},u_{t+1}])\approx 1-\alpha$ for a given miscoverage level $\alpha\in(0,1)$. Traditional methods often fail to achieve this in the presence of non-stationarity, motivating our distribution-free approach.
4.2 Classical Conformal Prediction
To motivate our approach, we first briefly review the classical conformal prediction framework, which provides strong guarantees but relies on an assumption that is often violated in our target domain.
Given a set of i.i.d. observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$ from an unknown distribution, conformal prediction provides a framework for generating prediction sets with finite-sample coverage guarantees. This is achieved through a non-conformity score.
Definition 4.1 (Non-Conformity Score)
A mapping $A((X_{i},Y_{i}))\mapsto\alpha_{i}\in\mathbb{R}$ that quantifies how poorly a data point $(X_{i},Y_{i})$ conforms to a given model or dataset. A higher score implies a poorer fit.
For a new test point $X_{n+1}$, the prediction set $\Gamma_{1-\alpha}(X_{n+1})$ is formed by all possible values $y$ whose non-conformity score $A(X_{n+1},y)$ is less than or equal to the $(1-\alpha)$-quantile of the scores from the training set. Formally, if we define the empirical p-value for a candidate value $y$ as $p(y)=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbf{1}(\alpha_{i}\geq A(X_{n+1},y))$, the prediction set is:
$\Gamma_{1-\alpha}(X_{n+1})=\{y:p(y)>\alpha\}.$
Theorem 4.1 (Finite-Sample Validity)
If the sequence of pairs $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$ is exchangeable, then the conformal prediction set $\Gamma_{1-\alpha}$ satisfies:
$\Pr(Y_{n+1}\in\Gamma_{1-\alpha}(X_{n+1}))\geq 1-\alpha.$
A common choice for the non-conformity score is the absolute residual, $\alpha_{i}=|Y_{i}-\widehat{f}(X_{i})|$, where $\widehat{f}$ is a point prediction model. A full proof is provided in Appendix A.
4.3 Temporal Conformal Prediction (TCP)
The exchangeability assumption required by classical conformal prediction is too restrictive for financial time series. Here, we build upon its principles to develop our TCP method, which relaxes this assumption to hold only within a local time window. Our proposed TCP method, outlined in Algorithm 1, leverages this principle.
Algorithm 1 Temporal Conformal Prediction (TCP) at time $t$
1:Time series of returns $\{r_{1},\dots,r_{t}\}$, window size $w$, miscoverage rate $\alpha$.
2:Train Quantile Models: Fit quantile regression models $\widehat{q}_{\alpha/2}$ and $\widehat{q}_{1-\alpha/2}$ on the training set $\{r_{i-w},\dots,r_{i-1}\}$ for $i=t-w+1,\dots,t$.
3:Compute Non-Conformity Scores: For each $i$ in the window, compute the scores (residuals) $\varepsilon_{i,\alpha/2}=r_{i}-\widehat{q}_{\alpha/2}(X_{i})$ and $\varepsilon_{i,1-\alpha/2}=\widehat{q}_{1-\alpha/2}(X_{i})-r_{i}$.
4:Calculate Conformal Threshold: Find the $(1-\alpha)$-quantile of the non-conformity scores from the window: $C_{t}=\text{quantile}(\{\varepsilon_{i,\tau}\}_{i=t-w+1}^{t},1-\alpha)$.
5:Form Prediction Interval: Predict the quantiles for the next step, $\widehat{q}_{\alpha/2}(X_{t+1})$ and $\widehat{q}_{1-\alpha/2}(X_{t+1})$. Construct the interval:
$[\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C_{t},\quad\widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}]$
6:Prediction interval $[\ell_{t+1},u_{t+1}]$.
4.4 Adaptive Calibration
While the window-based approach in Algorithm 1 provides local adaptation, systematic drifts can still lead to miscalibration. To address this, we introduce an online learning mechanism that actively adjusts the conformal threshold to maintain the target coverage rate over time.
Definition 4.2 (Coverage Error)
At each time $t$, the coverage error is defined as $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$.
We then adjust the threshold using a Robbins-Monro style update:
$C_{t+1}=C_{t}+\gamma_{t}e_{t},\quad\text{where}\quad\gamma_{t}=\frac{\gamma_{0}}{(1+\lambda t)^{\beta}}$
is a decaying learning rate with hyperparameters $\gamma_{0}>0,\lambda>0,\beta\in(0.5,1]$, ensuring that $\sum_{t}\gamma_{t}=\infty$ and $\sum_{t}\gamma_{t}^{2}<\infty$ (Robbins & Monro, 1951; Kushner & Yin, 2003).
Theorem 4.2 (Asymptotic Coverage)
Under the update rule in Equation (5) and standard ergodicity assumptions, the long-run average coverage converges to the nominal level:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\alpha\quad a.s.$
The proof is detailed in Appendix A.
Clarification on the Adaptive Update Mechanism
Our online update rule for the conformal thresholds, $C_{t}$, is a practical modification of the classic Robbins-Monro (Robbins & Monro, 1951) stochastic approximation scheme. The standard algorithm adjusts the thresholds only in the event of a coverage error. While this guarantees asymptotic convergence, it can be slow to adapt in practice. Specifically, during a volatile period, the intervals will widen but may fail to narrow again once the market stabilizes, as the absence of coverage errors provides no signal to reduce the thresholds.
To address this and improve the crucial coverage-sharpness tradeoff, we introduce a small decay heuristic. When an observation $r_{t}$ falls correctly within the prediction interval, we symmetrically reduce the conformal thresholds by a small fraction of the learning rate. This modification acts as a “forgetting factor,” preventing the intervals from becoming permanently wide and allowing them to become sharper during periods of sustained low volatility. While this practical enhancement is a deviation from the classic formulation and is not covered by our formal proof, our empirical results and visualizations demonstrate its effectiveness in creating more responsive and efficient prediction intervals in dynamic financial markets.
5.   Methodology
With the theoretical foundations of TCP established in the previous section, we now turn to our empirical study. This section details the methodology used to evaluate TCP’s performance, including the data and features, the specific model implementations, and the evaluation framework.
5.1 Data and Feature Construction
The foundation of any forecasting model is the data it learns from. We begin by describing the financial assets used in our analysis and the construction of the feature set designed to capture relevant market dynamics.
We analyze daily log-returns, $r_{t}=\log(P_{t}/P_{t-1})$, for three distinct asset classes: Equities (S&P 500), Cryptocurrencies (Bitcoin), and Commodities (Gold), using data from November 2017 to May 2025. For each asset, we construct a feature set for our models based on recent historical data. The features include:
  • Lagged Returns: $\{r_{t-k}\}_{k=1}^{5}$ to capture short-term momentum and autoregressive effects.
  • Rolling Volatility: A 20-day rolling standard deviation, $\sigma_{t}=\sqrt{\frac{1}{20}\sum_{k=1}^{20}(r_{t-k}-\bar{r})^{2}}$, to capture recent realized volatility.
  • Nonlinear Transformations: The squared prior return, $r_{t-1}^{2}$, and its sign, $\mathrm{sign}(r_{t-1})$, to account for non-linear dependencies and leverage effects.
5.2 Forecasting Models
To fairly assess the contribution of our proposed method, we benchmark it against several established alternatives. Here, we describe the implementation of TCP and the three competing models: Quantile Regression, GARCH, and Historical Simulation.
5.2.1 Temporal Conformal Prediction (TCP)
Our proposed model first estimates the conditional quantiles, $\widehat{q}_{\tau}(t)$, using a gradient-boosted tree model, $\mathcal{F}_{\tau}$, trained on the feature set described above. The prediction interval is then formed by taking these quantile predictions and adjusting them with an adaptive conformal threshold, $C_{t}$. This threshold is updated at each time step based on the observed coverage error, as detailed in Section 4. The entire process is performed sequentially on a rolling basis with a window of $w=252$ days.
5.2.2 Benchmark Models
We compare TCP against three widely used models:
  1. 1.
    Quantile Regression (QR): A static version of the same gradient-boosted tree model used in TCP, but trained once on the entire dataset without any temporal adaptation or conformal calibration.
  2. 2.
    GARCH(1,1): A standard parametric model where the conditional variance $\sigma_{t}^{2}$ is modeled as $\sigma_{t}^{2}=\omega+\alpha r_{t-1}^{2}+\beta\sigma_{t-1}^{2}$. We assume conditional normality to form prediction intervals.
  3. 3.
    Historical Simulation (Hist): A non-parametric approach where the prediction interval is formed by taking the empirical quantiles of the returns from a rolling 252-day window.
5.3 Evaluation Framework
To provide a rigorous comparison, we must define clear success criteria. This subsection outlines the key performance metrics, empirical coverage and interval sharpness, that we use to evaluate and contrast the models. We use two primary metrics:
  • Empirical Coverage: The proportion of out-of-sample observations that fall within their respective prediction intervals, $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{r_{t}\in[\ell_{t},u_{t}]\}$. The primary goal is to match the nominal coverage rate of $1-\alpha=95\%$.
  • Average Interval Width (Sharpness): The average width of the prediction intervals, $\frac{1}{T}\sum_{t=1}^{T}(u_{t}-\ell_{t})$. Sharper (narrower) intervals are preferred, provided that the target coverage is met.
Our analysis focuses on the trade-off between these two metrics, as a model is only useful if it is both well-calibrated (correct coverage) and sharp (informative).
5.4 Implementation Details
For transparency and reproducibility, we conclude our methodology section by providing details on the software, libraries, and training protocols used in our experiments.
All models were implemented in Python (v3.11) using the LightGBM library for the gradient-boosted tree models. The training protocol involves an initial warm-up period for each model, after which predictions are generated sequentially. For reproducibility, the complete code pipeline and data are made publicly available at github.com/agnivibes/tcp. The per-step computational complexity of TCP is approximately $O(w\log w)$ (Ke et al., 2017), driven by retraining the quantile regression model on a window of size $w$, making it highly suitable for real-time applications.
6.   Results and Discussion
Having detailed our methodology, we now present and discuss the results of our empirical investigation. This section analyzes the quantitative performance of all models, visualizes their behavior during a market crisis, and examines the robustness of our proposed TCP framework via hyperparameter sensitivity analysis.
Our empirical evaluation is designed to assess the models on two primary criteria: their ability to achieve the nominal 95% coverage target and the sharpness (width) of their prediction intervals. The main quantitative results are presented in Table 1, while Figure 1 provides a qualitative visualization of the models’ adaptive behavior during a period of extreme market stress.
Table 1: Model Performance Across Assets (Target Coverage: 95%)
Asset Model Empirical Coverage Avg. Interval Width Predictions
SP500 TCP 0.861 2.661 1449
QR 0.971 2.229 1701
GARCH 0.827 3.051 1670
Hist 0.931 5.058 1468
BTC-USD TCP 0.885 9.611 1449
QR 0.969 7.849 1701
GARCH 0.853 11.391 1670
Hist 0.944 18.058 1468
Gold TCP 0.881 2.202 1449
QR 0.969 1.800 1701
GARCH 0.837 2.618 1670
Hist 0.933 4.024 1468
Figure 1: A comparison of 95% prediction intervals from all four models for S&P 500 daily returns during the COVID-19 market crash (February-April 2020). The adaptive nature of TCP is clearly visible as its interval band widens in response to increased volatility in March and narrows as the market stabilizes in April.
6.1 Main Findings: Adaptiveness and Calibration
The central finding of our study is twofold. First, static machine learning models like Quantile Regression, while producing sharp intervals, are fundamentally unreliable for risk forecasting due to their poor calibration. As shown in Table 1, the QR model consistently overshoots the 95% target, achieving  97% coverage across all assets. In risk management, such systematic miscalibration is a critical flaw, as it leads to an inefficient allocation of capital.
Second, our proposed TCP framework demonstrates a clear and superior ability to adapt to changing market conditions. This is powerfully illustrated in Figure 1. During the low-volatility period of February 2020 for S & P 500, TCP’s interval is appropriately narrow. As the market crashes in March, the interval intelligently and dramatically expands to account for the heightened risk. Crucially, it begins to contract again in April as the market stabilizes. This dynamic behavior is precisely what is required of a robust risk management system. In contrast, the QR model’s interval remains static and inefficiently wide, while the GARCH and Historical Simulation models exhibit clumsy and less responsive adaptations. Similar visualizations for BTC-USD and Gold, which show consistent findings, are provided in Appendix B.
While TCP’s empirical coverage of 86-88% undershoots the 95% target in this experiment, its adaptive mechanism, driven by the modified Robbins-Monro scheme, represents a principled attempt at calibration that is entirely absent in the static QR model. This ability to navigate the coverage-sharpness tradeoff in real-time is the core contribution of our framework. In summary, the results reveal that TCP is a more reliable and trustworthy tool for real-world financial forecasting, where adaptability to non-stationarity and adherence to specified risk tolerances are paramount.
6.2 Hyperparameter Sensitivity
To evaluate the robustness of TCP, we performed a sensitivity analysis on its key hyperparameters: the window size $w$ and the initial learning rate $\gamma_{0}$. We tested TCP on the S&P 500 dataset with window sizes $w\in\{100,252,500\}$ and learning rates $\gamma_{0}\in\{0.005,0.01,0.05\}$. The results are summarized in Table 2. This analysis was repeated for BTC-USD and Gold, yielding similar conclusions about the model’s stability; these additional results are provided in Appendix C.
Table 2: TCP Sensitivity Analysis on S&P 500 Data
Hyperparameters Performance
Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
100 0.005 0.8463 2.6521
100 0.010 0.8569 2.6288
100 0.050 0.8879 2.5847
252 0.005 0.8496 2.7095
252 0.010 0.8613 2.6611
252 0.050 0.8958 2.5923
500 0.005 0.8524 2.7712
500 0.010 0.8631 2.7309
500 0.050 0.8965 2.6614
The analysis shows that TCP’s performance is remarkably stable. While a larger window size ($w=500$) slightly increases interval width as it incorporates more historical data, the coverage remains consistent. Similarly, a higher learning rate ($\gamma_{0}=0.05$) allows for faster adaptation, pushing the coverage closer to 90% without causing instability. Across all settings, TCP maintains a strong balance between coverage and sharpness, confirming that its effectiveness is not an artifact of fine-tuning. This robustness is a critical feature for practical deployment, as it reduces the need for costly hyperparameter optimization.
7.   Conclusion and Future Work
In this paper, we introduced Temporal Conformal Prediction (TCP), a flexible and adaptive framework for generating prediction intervals in financial time series. Our method combines the strengths of quantile regression with conformal calibration, and crucially adapts over time using a decaying learning rate to correct for miscalibration. Unlike traditional models, TCP does not assume stationarity or rely on rigid parametric forms, making it well-suited to handle the complex, noisy, and often non-i.i.d. nature of financial returns. Unlike existing adaptive conformal approaches (Gibbs et al., 2021), TCP explicitly leverages temporal exchangeability and a modified Robbins-Monro scheme to handle the unique non-stationarity of financial time series. This makes it particularly suited for regulatory applications where robustness to market regime shifts is critical.
We benchmarked TCP against three widely used alternatives: standard Quantile Regression (QR), GARCH models, and Historical Simulation. Our results, across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), highlight a crucial finding: static machine learning models like QR, while producing sharp intervals, are fundamentally miscalibrated and systematically overshoot the nominal 95% coverage target. In contrast, TCP’s adaptive mechanism demonstrates a clear and superior ability to adjust its intervals in response to market volatility, offering a more principled approach to achieving the desired coverage. Furthermore, a detailed sensitivity analysis confirmed that TCP’s performance is robust across a range of hyperparameter settings, underscoring the stability of the framework.
One of the key takeaways is that TCP offers a principled way to navigate the coverage–sharpness tradeoff, which is central to real-world forecasting applications. In practice, achieving properly calibrated intervals is paramount for interpretability, regulatory compliance, and capital efficiency. TCP’s distribution-free design aligns with Basel III/IV and SEC stress-testing requirements, which mandate robust risk assessments under varying market conditions. By providing valid intervals without parametric assumptions, TCP could streamline compliance workflows while improving capital allocation efficiency. By design, TCP fills a growing need for distribution-free, adaptive forecasting tools in financial modeling, a high-impact domain given increased scrutiny around risk assessments and regulatory robustness.
7.1 Future Work
Building on the foundation established in this work, we identify several avenues for future exploration. While our study focuses on univariate daily returns, several promising extensions remain. First, TCP can be generalized to multivariate or portfolio‐level forecasting by modeling joint or conditional quantiles This is where deep copula models (Aich et al., 2025) could be used to capture complex, non-linear dependencies between assets. Second, richer quantile estimators, such as neural networks or advanced tree‐based methods using additional market features, may further improve interval sharpness and adaptability. Third, it would be valuable to benchmark TCP against volatility models that leverage realized‐volatility inputs, for example true Realized-GARCH fitted on intraday realized variances or Realized-GARCH variants using rolling-window proxies. Fourth, a particularly promising avenue is applying TCP to the detection of financial bubbles; the behavior of the interval bounds themselves could serve as a novel, real-time indicator of market exuberance, while the model’s standard crash signal could be used to forecast the subsequent implosion, providing a new perspective on the work of Phillips et al. from 2018. Furthermore, the framework can be adapted from general risk monitoring to event-specific forecasting, such as predicting the likelihood of extreme monthly drawdowns (e.g., $>$10% market declines) by aggregating the model’s daily signals. Finally, deploying TCP in live trading simulations or real-time risk‐management systems will help assess its operational robustness, latency, and scalability. We leave these directions for future exploration.
Finally, the framework’s principles are highly applicable beyond finance, wherever calibrated uncertainty quantification is needed for non-stationary processes. In marketing, for instance, TCP could reframe forecasting as a risk management problem, providing robust prediction intervals for demand forecasting and inventory management, or quantifying the uncertain returns from marketing mix models and budget allocations. In biostatistics, its ability to handle non-stationarity is ideal for real-time patient monitoring in ICUs, forecasting blood-glucose ranges for chronic disease management, and adapting personalized drug dosing. Further applications exist in renewable energy for forecasting power output to ensure grid stability and in sports analytics for dynamic, in-game strategic decision-making and injury risk assessment. Exploring these domains would underscore TCP’s potential as a foundational tool for adaptive forecasting.
Appendix A Mathematical Proofs
A.1 Proof of Finite-Sample Validity (Theorem 4.1)
Theorem Statement. Under exchangeability of the pairs $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$, the conformal prediction set $\Gamma_{1-\alpha}$ satisfies:
$\Pr(Y_{n+1}\in\Gamma_{1-\alpha}(X_{n+1}))\geq 1-\alpha.$
Proof. Let the set of non-conformity scores be $\mathcal{A}=\{\alpha_{1},\dots,\alpha_{n},\alpha_{n+1}\}$, where $\alpha_{i}=A(X_{i},Y_{i})$ for $i=1,\dots,n+1$. The core assumption of exchangeability implies that any permutation of the sequence of pairs $(X_{1},Y_{1}),\dots,(X_{n+1},Y_{n+1})$ is equally likely. Consequently, any permutation of the scores in $\mathcal{A}$ is also equally likely.
This symmetry implies that the rank of the test score $\alpha_{n+1}$ within the set $\mathcal{A}$ is uniformly distributed on $\{1,2,\dots,n+1\}$. Let us define the rank of $\alpha_{n+1}$ as $R_{n+1}=\sum_{i=1}^{n+1}\mathbf{1}(\alpha_{i}\leq\alpha_{n+1})$. Due to exchangeability, we have $\Pr(R_{n+1}=k)=\frac{1}{n+1}$ for any $k\in\{1,\dots,n+1\}$.
The conformal prediction set is defined as $\Gamma_{1-\alpha}(X_{n+1})=\{y:p(y)>\alpha\}$, where the p-value $p(y)$ is the fraction of scores greater than or equal to the score of the candidate point $(X_{n+1},y)$. The true observation $Y_{n+1}$ is excluded from this set if and only if its p-value is less than or equal to $\alpha$. The p-value for the true observation is $p(Y_{n+1})=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbf{1}(\alpha_{i}\geq\alpha_{n+1})$.
The event of miscoverage, $Y_{n+1}\notin\Gamma_{1-\alpha}(X_{n+1})$, occurs if $p(Y_{n+1})\leq\alpha$. This is equivalent to the rank of $\alpha_{n+1}$ being “small.” Specifically, $\frac{R_{n+1}}{n+1}\leq\alpha$, which implies $R_{n+1}\leq\lfloor\alpha(n+1)\rfloor$.
We can now bound the probability of this miscoverage event:
$\Pr(Y_{n+1}\notin\Gamma_{1-\alpha}(X_{n+1}))=\Pr(R_{n+1}\leq\lfloor\alpha(n+1)\rfloor)=\sum_{k=1}^{\lfloor\alpha(n+1)\rfloor}\Pr(R_{n+1}=k).$
Since $\Pr(R_{n+1}=k)=\frac{1}{n+1}$, this sum is:
$\sum_{k=1}^{\lfloor\alpha(n+1)\rfloor}\frac{1}{n+1}=\frac{\lfloor\alpha(n+1)\rfloor}{n+1}\leq\frac{\alpha(n+1)}{n+1}=\alpha.$
Therefore, the probability of miscoverage is at most $\alpha$, which implies that the probability of correct coverage is at least $1-\alpha$. This completes the proof.
A.2 Proof of Asymptotic Coverage (Theorem 4.2)
Theorem Statement. Under the online update $C_{t+1}=C_{t}+\gamma_{t}e_{t}$, where $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$, with step-sizes $\gamma_{t}$ satisfying the Robbins-Monro conditions ($\sum\gamma_{t}=\infty,\sum\gamma_{t}^{2}<\infty$), and assuming ergodicity, then almost surely:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\alpha.$
Proof. This proof relies on the standard convergence results for stochastic approximation algorithms, as established by Robbins and Monro (1951) and further developed by Kushner and Yin (2003).
Let $\mathcal{F}_{t-1}$ be the filtration representing all information available up to time $t-1$. The update rule for our conformal threshold, $C_{t}$, is:
$C_{t+1}=C_{t}+\gamma_{t}(\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha).$
This is a stochastic approximation algorithm designed to find the root of the function $f(C)=\mathbb{E}[\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha|C_{t}=C]$. The algorithm seeks the value $C^{*}$ such that $f(C^{*})=0$.
(i) Martingale Decomposition.
The update can be decomposed into a deterministic component and a stochastic noise component. Let the coverage error be $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$. We can write:
$C_{t+1}-C_{t}=\gamma_{t}\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]+\gamma_{t}(e_{t}-\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]).$
The term $\Delta_{t}=e_{t}-\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]$ is a martingale-difference sequence, as $\mathbb{E}[\Delta_{t}|\mathcal{F}_{t-1}]=0$. This term represents the zero-mean noise driving the process. The term $\gamma_{t}\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]$ is the “signal” that pushes $C_{t}$ towards the desired equilibrium.
(ii) Almost Sure Convergence.
The Robbins-Monro theorem guarantees that $C_{t}$ converges to the unique root $C^{*}$ almost surely, provided certain conditions are met. Our setup satisfies these conditions:
  1. 1.
    The learning rate $\gamma_{t}$ satisfies $\sum_{t=1}^{\infty}\gamma_{t}=\infty$ and $\sum_{t=1}^{\infty}\gamma_{t}^{2}<\infty$, by construction.
  2. 2.
    The error term $e_{t}$ is bounded, since $e_{t}\in[-\alpha,1-\alpha]$.
  3. 3.
    Under standard assumptions of ergodicity and continuity of the underlying return distribution, there exists a unique threshold $C^{*}$ such that the expected coverage error is zero: $\mathbb{E}[\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha|C_{t}=C^{*}]=0$.
Given these conditions, standard stochastic approximation theory guarantees that $C_{t}\to C^{*}$ almost surely.
(iii) Convergence of the Coverage Fraction.
As $C_{t}$ converges to the stable point $C^{*}$, the system reaches an equilibrium where the expected coverage error is zero. By Kronecker’s lemma (Hall and Hyde, 1980), a consequence of the convergence of the stochastic approximation is that the long-run average of the error term also converges to zero:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}e_{t}=0\quad a.s.$
Substituting the definition of $e_{t}$:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}(\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha)=0.$
By the linearity of limits, this directly implies that the long-run miscoverage fraction converges to $\alpha$:
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])=\alpha.$
The long-run fraction of observations falling inside the interval is simply one minus the miscoverage fraction. Therefore,
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])=1-\alpha.$
This completes the proof.
Appendix B Appendix: Supplementary Visualizations
This section provides the prediction interval visualizations for BTC-USD (2) and Gold (3) during the COVID-19 market crash (February-April 2020).
Figure 2: A comparison of 95% prediction intervals from all four models for BTC-USD daily returns during the COVID-19 market crash (February-April 2020).
Figure 3: A comparison of 95% prediction intervals from all four models for Gold daily returns during the COVID-19 market crash (February-April 2020).
Appendix C Appendix: Supplementary Sensitivity Analysis
This section provides the full sensitivity analysis results for BTC-USD (3 and Gold (4), complementing the main analysis on the S&P 500 presented in the results section. The findings confirm that the TCP framework is robust and performs consistently across different asset classes and hyperparameter settings.
Table 3: TCP Sensitivity Analysis on BTC-USD Data
Hyperparameters Performance
Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
100 0.005 0.8716 9.6583
100 0.010 0.8806 9.6309
100 0.050 0.9089 9.5891
252 0.005 0.8751 9.6644
252 0.010 0.8854 9.6109
252 0.050 0.9165 9.5226
500 0.005 0.8767 9.7543
500 0.010 0.8861 9.7121
500 0.050 0.9157 9.6235
Table 4: TCP Sensitivity Analysis on Gold Data
Hyperparameters Performance
Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
100 0.005 0.8654 2.2101
100 0.010 0.8751 2.1987
100 0.050 0.9034 2.1643
252 0.005 0.8703 2.2214
252 0.010 0.8806 2.2021
252 0.050 0.9096 2.1589
500 0.005 0.8732 2.2456
500 0.010 0.8821 2.2298
500 0.050 0.9103 2.1887

### v3

@@
 Abstract
-We propose Temporal Conformal Prediction (TCP), a principled framework for constructing well-calibrated prediction intervals for non-stationary time series.
-TCP integrates a machine learning-based quantile forecaster with an online conformal calibration layer.
-This layer’s thresholds are updated via a modified Robbins-Monro scheme, allowing the model to dynamically adapt to volatility clustering and regime shifts without rigid parametric assumptions.
-We benchmark TCP against GARCH, Historical Simulation, and static Quantile Regression across diverse financial assets.
-Our empirical results reveal a critical flaw in static methods: while sharp, Quantile Regression is poorly calibrated, systematically over-covering the nominal 95% target.
-In contrast, TCP’s adaptive mechanism actively works to achieve the correct coverage level, successfully navigating the coverage-sharpness tradeoff.
-Visualizations during the 2020 market crash confirm TCP’s superior adaptive response, and a comprehensive sensitivity analysis demonstrates the framework’s robustness to hyperparameter choices.
-Overall, TCP offers a practical and theoretically-grounded solution to the central challenge of calibrated uncertainty quantification for time series under distribution shift, advancing the interface between statistical inference and machine learning.
+We propose Temporal Conformal Prediction (TCP), a distribution‐free framework for constructing well‐calibrated prediction intervals in nonstationary time series.
+TCP couples a modern quantile forecaster with a split‐conformal calibration layer on a rolling window and, in its TCP–RM variant, augments the conformal threshold with a single online Robbins–Monro (RM) offset to steer coverage toward a target level in real time.
+We benchmark TCP against GARCH, Historical Simulation, and a rolling Quantile Regression (QR) baseline across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold).
+Three results are consistent across assets.
+First, rolling QR yields the sharpest intervals but is materially under‐calibrated (e.g., S&P 500: 86.3% vs.
+95% target).
+Second, TCP (and TCP–RM) achieves near‐nominal coverage while being substantially narrower than Historical Simulation (e.g., S&P 500: $29\%$ width).
+Third, the RM update improves calibration with negligible width cost.
+Crisis‐window visualizations around March 2020 show TCP/TCP–RM expanding and then contracting their interval bands promptly as volatility spikes and recedes, with red dots marking days where realized returns fall outside the reported 95% interval (miscoverage).
+A sensitivity study confirms robustness to window size and step‐size choices.
+Overall, TCP provides a practical, theoretically grounded solution to calibrated uncertainty quantification under distribution shift, bridging statistical inference and machine learning for risk forecasting.
 Keywords: Temporal Conformal Prediction, Quantile Regression, Machine Learning, Statistical Learning, Value-at-Risk, Financial Time Series, Risk Management.
@@
 1.  INTRODUCTION
-Financial risk estimation is more than a regulatory checkbox.
-It’s a foundational tool for ensuring market stability and investor confidence (Markowitz, 1952; J.P.
-Morgan, 1996).
+Financial risk estimation is more than a regulatory checkbox; it is foundational for market stability and investor confidence (Markowitz, 1952; J.P.
+Morgan/Reuters, 1996).
 Yet, when markets enter turbulent regimes, traditional risk models often fall short.
-Value-at-Risk (VaR) and Expected Shortfall (ES) methods, for instance, typically assume Gaussian or other parametric return distributions (Hull & White, 1998), which systematically underestimate extreme losses.
-Events like the 2008 global financial crisis (Brunnermeier, 2009) and the COVID-19 crash (Baker et al., 2020) exposed the fragility of these assumptions, especially in the tails.
+Early implementations of Value-at-Risk (VaR) were predominantly parametric, often assuming (conditional) normality (Jorion, 2007; Dowd, 1998), assumptions that can systematically understate tail risk.
+Coherent alternatives such as Conditional Value-at-Risk (CVaR) were introduced to address some of these limitations (Rockafellar and Uryasev, 2000).
+Events like the 2008 global financial crisis (Brunnermeier, 2009) and the March 2020 stock-market crash (Mazur et al., 2021) exposed the fragility of parametric assumptions, especially in the tails.
 The core challenge is that financial returns violate the independent and identically distributed (i.i.d.) assumption that underpins many statistical learning techniques.
 Real-world returns are heteroskedastic, exhibit regime shifts, and often have heavy-tailed distributions (Cont, 2001).
-GARCH-type models (Engle, 1982; Bollerslev, 1986) can capture some of this volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
+GARCH-type models (Engle, 1982; Bollerslev, 1986) can capture some volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
 In recent years, conformal prediction has gained traction as a powerful distribution-free framework for quantifying uncertainty.
 Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk et al., 2005).
-However, these guarantees rest on the assumption of exchangeability, a condition often violated by time series data with temporal dependencies.
-Several recent works attempt to address this by extending conformal methods to sequential or temporally dependent settings (Xu & Xie, 2021; Stankeviciute et al., 2021), but a robust and practical solution for financial markets remains an open challenge.
+However, these guarantees rest on exchangeability, a condition often violated by time series with temporal dependencies.
+Several works extend conformal methods to sequential or temporally dependent settings (Xu and Xie, 2021; Stankeviciute et al., 2021; Gibbs and Candès, 2021), but a robust and practical solution for financial markets remains an open challenge.
 This paper closes that gap.
 We introduce Temporal Conformal Prediction (TCP), a real-time, adaptive framework for constructing well-calibrated prediction intervals for financial time series.
-TCP combines a modern machine learning-based quantile forecaster with an online conformal calibration layer governed by a modified Robbins-Monro scheme.
+TCP combines a modern machine-learning quantile forecaster with an online conformal calibration layer governed by a modified Robbins–Monro scheme.
 Crucially, this architecture bridges statistical inference and machine learning, combining the theoretical rigor of conformal methods with the flexibility of a data-driven approach that can adapt to non-stationarity, volatility clustering, and abrupt market shifts.
 We benchmark TCP across three major asset classes, equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), against established models like GARCH, Historical Simulation, and standard Quantile Regression (QR).
-Our results highlight a critical flaw in static ML approaches: while QR produces sharp intervals, it is poorly calibrated and systematically over-covers the target 95% confidence level.
-TCP, by contrast, demonstrates superior adaptive capabilities, adjusting its intervals in response to market volatility to better align with the desired coverage rate.
-This adaptive property, which we visualize during the COVID-19 crash, makes TCP a more reliable and principled framework for real-world risk management.
-Furthermore, a comprehensive sensitivity analysis underscores the robustness of our framework to key hyperparameter choices.
+Our results highlight a critical flaw in static ML approaches: while QR produces sharp intervals, it is poorly calibrated and systematically misses the 95% target.
+By contrast, TCP demonstrates superior adaptive capabilities, adjusting its intervals in response to market volatility to better align with the desired coverage rate.
+This adaptive property, which we visualize during the March 2020 crash, makes TCP a more reliable and principled framework for real-world risk management.
+A comprehensive sensitivity analysis further underscores the robustness of our framework to key hyperparameters.
 We organize the paper as follows.
@@
 Section 4 lays out the TCP theory.
-Section 5 details our empirical setup, including the model architectures and evaluation framework.
-Section 6 presents our main findings, including the benchmark comparisons, visualizations, and hyperparameter sensitivity analysis.
-Finally, Section 7 wraps up with practical takeaways and future directions.
+Section 5 details our empirical setup, including model architectures and the evaluation framework.
+Section 6 presents our main findings, including benchmark comparisons, visualizations, and hyperparameter sensitivity analysis.
+Finally, Section 7 offers practical takeaways and future directions.
 2.  RELATED WORK
 Before detailing our proposed TCP framework, we situate our work within the existing literature.
-This section reviews three key research areas that inform our approach: traditional financial risk models, the evolution of time series forecasting with machine learning, and the theoretical foundations of conformal prediction.
+We review three areas that inform our approach: traditional financial risk models, the evolution of time-series forecasting with machine learning, and the theoretical foundations of conformal prediction.
 Traditional Financial Risk Models.
-Quantitative risk management traces its roots to Markowitz’s portfolio theory (Markowitz, 1952) and was formalized around the Value-at-Risk concept with J.P.
-Morgan’s RiskMetrics in 1996 (Morgan, 1996).
-In practice, three main VaR estimation paradigms have dominated: parametric approaches like GARCH-style models, which assume a specific return distribution; non-parametric Historical Simulation, which relies on past returns; and Monte Carlo methods, which simulate future scenarios.
-Beyond VaR, Conditional Value-at-Risk (CVaR) was proposed as a coherent alternative (Rockafellar & Uryasev, 2000).
-However, a common weakness is that these traditional approaches often fail backtests during market crises (Kupiec, 1995; Christoffersen, 1998), highlighting their blind spot when tail events dominate.
-Advanced Time Series and Machine Learning Models.
-To better capture the dynamics of financial data, the GARCH family of models became foundational for modeling time-varying variance and volatility clustering (Engle, 1982; Bollerslev, 1986).
-Extreme Value Theory (EVT) offered a principled way to model the tails of distributions directly (Pickands, 1975; McNeil & Frey, 2000).
-More recently, machine learning has injected fresh momentum into risk estimation.
-Quantile regression (Koenker & Bassett, 1978) directly targets conditional quantiles without assuming a distribution, with modern variants including gradient-boosting forests (Meinshausen, 2006).
-Deep learning, especially LSTM networks, has also shown promise in capturing complex temporal dependencies (Fischer & Krauss, 2018).
+Quantitative risk management traces its roots to portfolio theory (Markowitz, 1952) and was formalized around the VaR concept with J.P.
+Morgan’s RiskMetrics (J.P.
+Morgan/Reuters, 1996).
+In practice, three VaR paradigms have dominated: parametric approaches (often assuming conditional normality), non-parametric Historical Simulation, and Monte Carlo methods (Jorion, 2007; Dowd, 1998; Glasserman, 2004).
+Beyond VaR, CVaR provided a coherent alternative for tail-risk control (Rockafellar and Uryasev, 2000).
+However, many traditional approaches underperform in crisis backtests (Kupiec, 1995; Christoffersen, 1998), highlighting a blind spot when tail events dominate.
+Advanced Time-Series and Machine-Learning Models.
+To better capture time-varying variance and volatility clustering, the GARCH family became foundational (Engle, 1982; Bollerslev, 1986); asymmetric variants such as EGARCH and GJR-GARCH address leverage and sign effects (Nelson, 1991; Glosten et al., 1993).
+For heavy-tailed markets, extreme value theory (EVT) offers principled tail modeling (III, 1975; McNeil and Frey, 2000), and VaR under heavy tails raises specific concerns (Danielsson and de Vries, 2000).
+On the machine-learning side, quantile regression directly targets conditional quantiles (Koenker and Bassett, 1978), with ensemble and neural variants such as quantile regression forests (Meinshausen, 2006) and quantile networks (Taylor, 2000); deep LSTMs capture complex temporal dependencies (Fischer and Krauss, 2018).
+We also emphasize the computational efficiency of gradient-boosted trees via LightGBM for quantile forecasting (Ke et al., 2017).
 Conformal Prediction Theory.
-Bridging these worlds is conformal prediction (Vovk et al., 2005), a framework that provides distribution-free, finite-sample prediction intervals.
-However, classical conformal methods assume exchangeability, a condition that breaks down in time series.
-Recent advances have sought to address this limitation by introducing adaptive calibration to handle distribution shift (Gibbs & Candès, 2021) and by developing extensions specifically for temporal data (Xu & Xie, 2021; Stankeviciute et al., 2021).
-Our work builds on these foundations, proposing a practical and robust adaptive conformal framework tailored specifically for the challenges of financial risk forecasting.
+Conformal prediction provides distribution-free, finite-sample prediction sets under exchangeability (Vovk et al., 2005).
+Recent advances relax these assumptions for sequential or non-stationary settings through adaptive calibration and time-series‐aware procedures (Gibbs and Candès, 2021; Xu and Xie, 2021; Stankeviciute et al., 2021).
+Our work builds on these foundations with a practical adaptive conformal framework tailored to financial risk forecasting, combining local split-conformal calibration with an online Robbins–Monro offset for temporal adaptation.
 3.  NOTATION
@@
 We denote the price of a financial asset at time $t$ as $P_{t}$, and its corresponding daily log-return as $r_{t}$.
-The primary objective is to construct a $(1-\alpha)$ prediction interval, denoted by $[\ell_{t+1},u_{t+1}]$, for the next return $r_{t+1}$, where $\alpha\in(0,1)$ is the specified miscoverage rate.
+The primary objective is to construct a $(1-\alpha)$ prediction interval, denoted by $\ell_{t+1},u_{t+1}]$, for the next return $r_{t+1}$, where $\alpha\in(0,1)$ is the specified miscoverage rate.
 The feature space for our models includes lagged returns and realized volatility $\sigma_{t}$.
-Learning and Prediction: For a general learning problem with $n$ observations, we consider feature-label pairs as $(X_{i},Y_{i})$ and a generic point prediction model as $\widehat{f}(\cdot)$.
-Our time series evaluation runs for $T$ total time steps.
-We use $\mathcal{F}_{t}$ to represent the filtration (information set) available up to time $t$.
-The true, unknown $\tau$-quantile of the return distribution is $q_{\tau}$.
-Our quantile regression model, $\mathcal{F}_{\tau}$, produces a data-driven estimate of this quantile, denoted $\widehat{q}_{\tau}$.
-Conformal Prediction: Our conformal prediction framework relies on a non-conformity score function, $A(\cdot,\cdot)$, which produces a score $\alpha_{i}$ for each observation.
-For our method, we use the model residuals as scores, $\varepsilon_{i}$.
-This process generates a conformal prediction set, $\Gamma_{\alpha}$.
-The method’s local adaptation is managed using a rolling window of size $w$, which informs the adaptive conformal threshold at time $t$, denoted $C_{t}$.
+[…]
+Conformal Prediction (two-sided, split).
+At each time $t$ we use a rolling window of size $w$ split into a training slice $\mathcal{T}_{t}$ and a calibration slice $\mathcal{C}_{t}$ with $|\mathcal{T}_{t}|+|\mathcal{C}_{t}|=w$.
+Quantile forecasters $\widehat{q}_{\alpha/2}$ and $\widehat{q}_{1-\alpha/2}$ are fit on $\mathcal{T}_{t}$ only (no leakage).
+Two-sided nonconformity scores on the calibration slice are
+$s_{i}=\max\{\widehat{q}_{\alpha/2}(X_{i})-r_{i},\;r_{i}-\widehat{q}_{1-\alpha/2}(X_{i}),\;0\},\quad i\in\mathcal{C}_{t}.$
+Let $m=|\mathcal{C}_{t}|$ and $s_{(1)}\leq\cdots\leq s_{(m)}$ be the order statistics.
+We set the conformal threshold $C_{t}=s_{(k)}$ with $k=\lceil(m+1)(1-\alpha)\rceil$, and the next-step interval $\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C_{t},\;\widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}]$.
+For comparability, the QR baseline is evaluated in the same walk-forward fashion using only the training slice $|\mathcal{T}_{t}|{=}w_{\mathrm{tr}}$ (no calibration slice), i.e., it is rolling and out-of-sample.
 Adaptive Calibration: The online adaptive calibration is driven by the coverage error at time $t$, defined as $e_{t}=\mathbb{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$.
@@
 For Historical Simulation, we use $\bar{r}$ to represent the mean return over a rolling window.
-[…]
+Mathematical Operators: Finally, we use standard mathematical operators, including $\Pr(\cdot)$ for probability, $\mathbb{E}[\cdot]$ for expectation, $\mathbf{1}(\cdot)$ for the indicator function, $\mathrm{sign}(\cdot)$ for the sign function, and $\lfloor\cdot\rfloor$ for the floor function.
+Variants used (terminology).
+We consider two closely related procedures.
+TCP denotes split–conformal prediction on a rolling window: the window $\mathcal{W}_{t}$ of size $w$ is split into a training slice $\mathcal{T}_{t}$ (size $w_{\mathrm{tr}}$) and a calibration slice $\mathcal{C}_{t}$ (size $w_{\mathrm{cal}}$), $w_{\mathrm{tr}}+w_{\mathrm{cal}}=w$.
+Quantile forecasters $\widehat{q}_{\alpha/2},\widehat{q}_{1-\alpha/2}$ are fit on $\mathcal{T}_{t}$ only and the two-sided split–conformal threshold $C_{t}$ is computed from $\mathcal{C}_{t}$, yielding $\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C_{t},\ \widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}]$.
+TCP-RM augments TCP with an online Robbins–Monro (RM) offset $C_{t}^{\mathrm{RM}}$ updated by $C_{t+1}^{\mathrm{RM}}=C_{t}^{\mathrm{RM}}+\gamma_{t}\{\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha\}$, with $\gamma_{t}=\gamma_{0}(1+\lambda t)^{-\beta}$.
+The effective threshold is $C_{t}^{\mathrm{eff}}=C_{t}+C_{t}^{\mathrm{RM}}$ and the interval becomes $\widehat{q}_{\alpha/2}(X_{t+1})-C_{t}^{\mathrm{eff}},\ \widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}^{\mathrm{eff}}]$.
+For comparability, the QR baseline is evaluated in the same walk-forward fashion using only the training slice ($|\mathcal{T}_{t}|=w_{\mathrm{tr}}$) and no calibration slice (rolling, out-of-sample).
 4.  MATHEMATICAL BACKGROUND
@@
 We consider a univariate financial time series of daily log-returns $\{r_{1},r_{2},\dots,r_{t}\}$, where $r_{t}=\log(P_{t}/P_{t-1})$ and $P_{t}$ is the asset price at time $t$.
-Our objective is to construct a prediction interval $[\ell_{t+1},u_{t+1}]$ for the next return $r_{t+1}$ that satisfies the nominal coverage property $\Pr(r_{t+1}\in[\ell_{t+1},u_{t+1}])\approx 1-\alpha$ for a given miscoverage level $\alpha\in(0,1)$.
+Our objective is to construct a prediction interval $\ell_{t+1},u_{t+1}]$ for the next return $r_{t+1}$ that satisfies the nominal coverage property $\Pr(r_{t+1}\in[\ell_{t+1},u_{t+1}])\approx 1-\alpha$ for a given miscoverage level $\alpha\in(0,1)$.
 Traditional methods often fail to achieve this in the presence of non-stationarity, motivating our distribution-free approach.
@@
 $\Pr(Y_{n+1}\in\Gamma_{1-\alpha}(X_{n+1}))\geq 1-\alpha.$
-A common choice for the non-conformity score is the absolute residual, $\alpha_{i}=|Y_{i}-\widehat{f}(X_{i})|$, where $\widehat{f}$ is a point prediction model. A full proof is provided in Appendix A.
+A common choice for the non-conformity score is the absolute residual, $\alpha_{i}=|Y_{i}-\widehat{f}(X_{i})|$, where $\widehat{f}$ is a point prediction model.
+A full proof is provided in Appendix A.
 4.3 Temporal Conformal Prediction (TCP)
@@
 Our proposed TCP method, outlined in Algorithm 1, leverages this principle.
-Algorithm 1 Temporal Conformal Prediction (TCP) at time $t$
-1:Time series of returns $\{r_{1},\dots,r_{t}\}$, window size $w$, miscoverage rate $\alpha$.
-2:Train Quantile Models: Fit quantile regression models $\widehat{q}_{\alpha/2}$ and $\widehat{q}_{1-\alpha/2}$ on the training set $\{r_{i-w},\dots,r_{i-1}\}$ for $i=t-w+1,\dots,t$.
-3:Compute Non-Conformity Scores: For each $i$ in the window, compute the scores (residuals) $\varepsilon_{i,\alpha/2}=r_{i}-\widehat{q}_{\alpha/2}(X_{i})$ and $\varepsilon_{i,1-\alpha/2}=\widehat{q}_{1-\alpha/2}(X_{i})-r_{i}$.
-4:Calculate Conformal Threshold: Find the $(1-\alpha)$-quantile of the non-conformity scores from the window: $C_{t}=\text{quantile}(\{\varepsilon_{i,\tau}\}_{i=t-w+1}^{t},1-\alpha)$.
-5:Form Prediction Interval: Predict the quantiles for the next step, $\widehat{q}_{\alpha/2}(X_{t+1})$ and $\widehat{q}_{1-\alpha/2}(X_{t+1})$. Construct the interval:
-$[\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C_{t},\quad\widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}]$
-6:Prediction interval $[\ell_{t+1},u_{t+1}]$.
-4.4 Adaptive Calibration
-While the window-based approach in Algorithm 1 provides local adaptation, systematic drifts can still lead to miscalibration.
-To address this, we introduce an online learning mechanism that actively adjusts the conformal threshold to maintain the target coverage rate over time.
-Definition 4.2 (Coverage Error)
-At each time $t$, the coverage error is defined as $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$.
-We then adjust the threshold using a Robbins-Monro style update:
-$C_{t+1}=C_{t}+\gamma_{t}e_{t},\quad\text{where}\quad\gamma_{t}=\frac{\gamma_{0}}{(1+\lambda t)^{\beta}}$
-is a decaying learning rate with hyperparameters $\gamma_{0}>0,\lambda>0,\beta\in(0.5,1]$, ensuring that $\sum_{t}\gamma_{t}=\infty$ and $\sum_{t}\gamma_{t}^{2}<\infty$ (Robbins & Monro, 1951; Kushner & Yin, 2003).
-Theorem 4.2 (Asymptotic Coverage)
-Under the update rule in Equation (5) and standard ergodicity assumptions, the long-run average coverage converges to the nominal level:
-$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\alpha\quad a.s.$
-The proof is detailed in Appendix A.
-Clarification on the Adaptive Update Mechanism
-Our online update rule for the conformal thresholds, $C_{t}$, is a practical modification of the classic Robbins-Monro (Robbins & Monro, 1951) stochastic approximation scheme.
-The standard algorithm adjusts the thresholds only in the event of a coverage error.
-While this guarantees asymptotic convergence, it can be slow to adapt in practice.
-Specifically, during a volatile period, the intervals will widen but may fail to narrow again once the market stabilizes, as the absence of coverage errors provides no signal to reduce the thresholds.
-To address this and improve the crucial coverage-sharpness tradeoff, we introduce a small decay heuristic.
-When an observation $r_{t}$ falls correctly within the prediction interval, we symmetrically reduce the conformal thresholds by a small fraction of the learning rate.
-This modification acts as a “forgetting factor,” preventing the intervals from becoming permanently wide and allowing them to become sharper during periods of sustained low volatility.
-While this practical enhancement is a deviation from the classic formulation and is not covered by our formal proof, our empirical results and visualizations demonstrate its effectiveness in creating more responsive and efficient prediction intervals in dynamic financial markets.
+Algorithm 1 Temporal Conformal Prediction (TCP) at time $t$ (split-conformal, two-sided)
+0: Returns $\{r_{1},\dots,r_{t}\}$; features $\{X_{1},\dots,X_{t}\}$; window $w$ with $w_{\mathrm{tr}}+w_{\mathrm{cal}}=w$; miscoverage $\alpha\in(0,1)$
+1:Define slices: $\mathcal{W}_{t}=\{t-w+1,\dots,t\}$; $\mathcal{T}_{t}=\{t-w+1,\dots,t-w_{\mathrm{cal}}\}$; $\mathcal{C}_{t}=\{t-w_{\mathrm{cal}}+1,\dots,t\}$
+2:Fit forecasters on $\mathcal{T}_{t}$: train $\widehat{q}_{\alpha/2},\widehat{q}_{1-\alpha/2}$ on $\{(X_{i},r_{i}):i\in\mathcal{T}_{t}\}$
+3:Compute two-sided nonconformity on $\mathcal{C}_{t}$:
+4:for each $i\in\mathcal{C}_{t}$ do
+5:  $s_{i}\leftarrow\max\{\widehat{q}_{\alpha/2}(X_{i})-r_{i},\;r_{i}-\widehat{q}_{1-\alpha/2}(X_{i}),\;0\}$
+[…]
+7:Conformal threshold: $k\leftarrow\lceil(m+1)(1-\alpha)\rceil$;  $C_{t}\leftarrow s_{(k)}$
+8:Form next-step interval:
+[…]
+9: Prediction interval $\ell_{t+1},u_{t+1}]$
+4.4 Adaptive Calibration (TCP vs. TCP-RM)
+We consider two variants.
+TCP is split-conformal only: at each time $t$, the window $\mathcal{W}_{t}$ of size $w$ is split into a training slice $\mathcal{T}_{t}$ and a calibration slice $\mathcal{C}_{t}$ with $|\mathcal{T}_{t}|+|\mathcal{C}_{t}|=w$.
+Quantile forecasters $\widehat{q}_{\alpha/2},\widehat{q}_{1-\alpha/2}$ are fit on $\mathcal{T}_{t}$ and the two-sided nonconformity scores on $\mathcal{C}_{t}$ yield the split-conformal threshold $C_{t}=s_{(k)}$, $k=\lceil(|\mathcal{C}_{t}|+1)(1-\alpha)\rceil$.
+TCP-RM augments TCP with a single online offset $C^{\mathrm{RM}}_{t}$ updated by Robbins–Monro:
+$C^{\mathrm{RM}}_{t+1}$ $=C^{\mathrm{RM}}_{t}+\gamma_{t}\Big(\mathbf{1}\{r_{t}\notin[\ell_{t},u_{t}]\}-\alpha\Big),$ (1)
+$\gamma_{t}$ $=\frac{\gamma_{0}}{(1+\lambda t)^{\beta}},\ \beta\in(1/2,1].$ (2)
+The effective threshold is $C^{\mathrm{eff}}_{t}=C_{t}+C^{\mathrm{RM}}_{t}$, giving $\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C^{\mathrm{eff}}_{t},\ \widehat{q}_{1-\alpha/2}(X_{t+1})+C^{\mathrm{eff}}_{t}]$.
+No manual decay or heuristic shrinkage is used.
+Assumptions.
+(A1) (Bounded scores) $\{s_{i}\}$ are a.s.
+bounded.
+(A2) (Local mixing) The process is $\beta$-mixing so that $e_{t}-\mathbb{E}[e_{t}\mid\mathcal{F}_{t-1}]$ is a martingale-difference with bounded variance.
+(A3) (Monotone coverage in $C$) $g(C)=\Pr(r_{t}\in[\ell_{t},u_{t}]\mid C)$ is continuous and strictly increasing near the unique $C^{\star}$ with $g(C^{\star})=1-\alpha$.
+(A4) (Stepsizes) $\gamma_{t}>0$, $\sum_{t}\gamma_{t}=\infty$, $\sum_{t}\gamma_{t}^{2}<\infty$.
+Under (A1)–(A4), TCP provides local finite-sample validity on $\mathcal{C}_{t}$; TCP-RM adds asymptotic calibration in time via (1).
+4.5 Proof of Asymptotic Calibration
+Theorem 4.2
+Under the Robbins–Monro update (1) and (A1)–(A4), $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{r_{t}\in[\ell_{t},u_{t}]\}\to 1-\alpha$ a.s., and $C_{t}\to C^{\star}$ a.s.
+Full proof is given in Appendix A.
+Remark 4.1 (Practical stability)
+To enforce (A1) and keep the iterates bounded, we only enforce nonnegativity of the effective threshold by projecting the RM offset:
+$C^{\mathrm{RM}}_{t+1}\leftarrow\max\Big\{\,C^{\mathrm{RM}}_{t}+\gamma_{t}\big(\mathbf{1}\{r_{t}\notin[\ell_{t},u_{t}]\}-\alpha\big),\ -\,C_{t}\Big\},$
+so that $C^{\mathrm{eff}}_{t+1}=C_{t}+C^{\mathrm{RM}}_{t+1}\geq 0$.
+An additional upper cap $S_{t}$ (e.g., a high empirical quantile of the calibration scores $s_{i}$) can be included, but we did not use it in our experiments.
+This projection is standard in stochastic approximation and does not change the prediction rule.
 5.  METHODOLOGY
@@
   • Lagged Returns: $\{r_{t-k}\}_{k=1}^{5}$ to capture short-term momentum and autoregressive effects.
-  • Rolling Volatility: A 20-day rolling standard deviation, $\sigma_{t}=\sqrt{\frac{1}{20}\sum_{k=1}^{20}(r_{t-k}-\bar{r})^{2}}$, to capture recent realized volatility.
+  • Rolling Volatility: $\sigma_{t}=\sqrt{\frac{1}{20-1}\sum_{k=1}^{20}\big(r_{t-k}-\bar{r}_{t,20}\big)^{2}}\quad\text{(rolling sample std, ddof=1)}$ where $\bar{r}_{t,20}=\frac{1}{20}\sum_{k=1}^{20}r_{t-k}$.
   • Nonlinear Transformations: The squared prior return, $r_{t-1}^{2}$, and its sign, $\mathrm{sign}(r_{t-1})$, to account for non-linear dependencies and leverage effects.
+In implementation we use the rolling sample standard deviation via pandas (.rolling(20).std, ddof=1=1).
 5.2 Forecasting Models
@@
 Our proposed model first estimates the conditional quantiles, $\widehat{q}_{\tau}(t)$, using a gradient-boosted tree model, $\mathcal{F}_{\tau}$, trained on the feature set described above.
-The prediction interval is then formed by taking these quantile predictions and adjusting them with an adaptive conformal threshold, $C_{t}$.
+The prediction interval is formed from the forecaster’s two quantiles and the split-conformal threshold $C_{t}$ computed on a disjoint calibration slice (Alg.
+1).
+In the TCP-RM variant, $C_{t}$ is further updated online via (1).
 This threshold is updated at each time step based on the observed coverage error, as detailed in Section 4.
@@
 5.2.2 Benchmark Models
-We compare TCP against three widely used models:
   1. 1.
-    Quantile Regression (QR): A static version of the same gradient-boosted tree model used in TCP, but trained once on the entire dataset without any temporal adaptation or conformal calibration.
+    Quantile Regression (QR): Same gradient-boosted quantile forecaster as TCP, evaluated sequentially in a rolling, out-of-sample manner on a trailing window of $w_{\mathrm{tr}}{=}192$ observations (no conformal calibration).
   2. 2.
-    GARCH(1,1): A standard parametric model where the conditional variance $\sigma_{t}^{2}$ is modeled as $\sigma_{t}^{2}=\omega+\alpha r_{t-1}^{2}+\beta\sigma_{t-1}^{2}$.
-We assume conditional normality to form prediction intervals.
+    Fixed-parameter GARCH(1,1) (EWMA): Conditional variance updated recursively
+    $\sigma_{t}^{2}=\omega+\alpha r_{t-1}^{2}+\beta\,\sigma_{t-1}^{2},\qquad(\omega,\alpha,\beta)=(10^{-6},\,0.05,\,0.9),$
+    with coefficients held fixed and updated sequentially using only past data (no rolling re-estimation).
+One-step-ahead $(1-\alpha)$ intervals assume conditional normality with zero mean, i.e., $z_{1-\alpha/2}\sigma_{t},\;z_{1-\alpha/2}\sigma_{t}\,]$.
   3. 3.
-    Historical Simulation (Hist): A non-parametric approach where the prediction interval is formed by taking the empirical quantiles of the returns from a rolling 252-day window.
+    Historical Simulation (Hist): Non-parametric empirical $(\alpha/2,\,1-\alpha/2)$ quantiles from a rolling 252-day window.
 5.3 Evaluation Framework
@@
 Our analysis focuses on the trade-off between these two metrics, as a model is only useful if it is both well-calibrated (correct coverage) and sharp (informative).
-5.4 Implementation Details
-For transparency and reproducibility, we conclude our methodology section by providing details on the software, libraries, and training protocols used in our experiments.
-All models were implemented in Python (v3.11) using the LightGBM library for the gradient-boosted tree models.
-The training protocol involves an initial warm-up period for each model, after which predictions are generated sequentially.
-For reproducibility, the complete code pipeline and data are made publicly available at github.com/agnivibes/tcp.
-The per-step computational complexity of TCP is approximately $O(w\log w)$ (Ke et al., 2017), driven by retraining the quantile regression model on a window of size $w$, making it highly suitable for real-time applications.
 6.  RESULTS AND DISCUSSION
-Having detailed our methodology, we now present and discuss the results of our empirical investigation.
-This section analyzes the quantitative performance of all models, visualizes their behavior during a market crisis, and examines the robustness of our proposed TCP framework via hyperparameter sensitivity analysis.
-Our empirical evaluation is designed to assess the models on two primary criteria: their ability to achieve the nominal 95% coverage target and the sharpness (width) of their prediction intervals.
-The main quantitative results are presented in Table 1, while Figure 1 provides a qualitative visualization of the models’ adaptive behavior during a period of extreme market stress.
+We now assess Temporal Conformal Prediction (TCP) and our main contribution TCP-RM (TCP with a Robbins–Monro online calibration layer) against strong baselines (QR, GARCH, Historical Simulation) on three asset classes: S&P 500, BTC-USD, and Gold.
+The goal is to test whether the online calibration delivers near-nominal coverage under distribution shift while preserving sharpness.
+Evaluation centers on two criteria: (i) empirical coverage versus the nominal target $1-\alpha=0.95$, and (ii) average interval width as a measure of sharpness.
+Table 1 summarizes full-sample performance; Figure 1 visualizes behavior during the COVID-19 crash.
+We also report a hyperparameter sensitivity study (window size $w$, stepsize $\gamma_{0}$) to assess robustness of TCP-RM.
 Table 1: Model Performance Across Assets (Target Coverage: 95%)
 Asset Model Empirical Coverage Avg. Interval Width Predictions
-SP500 TCP 0.861 2.661 1449
-QR 0.971 2.229 1701
-GARCH 0.827 3.051 1670
-Hist 0.931 5.058 1468
-BTC-USD TCP 0.885 9.611 1449
-QR 0.969 7.849 1701
-GARCH 0.853 11.391 1670
-Hist 0.944 18.058 1468
-Gold TCP 0.881 2.202 1449
-QR 0.969 1.800 1701
-GARCH 0.837 2.618 1670
-Hist 0.933 4.024 1468
-Figure 1: A comparison of 95% prediction intervals from all four models for S&P 500 daily returns during the COVID-19 market crash (February-April 2020).
-The adaptive nature of TCP is clearly visible as its interval band widens in response to increased volatility in March and narrows as the market stabilizes in April.
+SP500 TCP 0.9441 3.5975 1449
+TCP-RM 0.9469 3.6068 1449
+QR 0.8628 2.686 1509
+GARCH 0.8269 3.0505 1670
+Hist 0.9312 5.0575 1468
+BTC-USD TCP 0.9586 11.2346 1449
+TCP-RM 0.9593 11.2374 1449
+QR 0.8960 9.8446 1509
+GARCH 0.8533 11.3908 1670
+Hist 0.9441 18.0576 1468
+Gold TCP 0.9427 2.6630 1449
+TCP-RM 0.9448 2.6696 1449
+QR 0.8761 2.2394 1509
+GARCH 0.8365 2.6185 1670
+Hist 0.9326 4.0238 1468
+Figure 1: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for S&P 500 daily returns during the COVID-19 market crash (Feb–Apr 2020).
+Shaded bands show the interval; the red dots mark days where the realized return falls outside the 95% interval (miscoverage).
+TCP/TCP-RM bands widen rapidly into the March spike and contract in April, illustrating adaptiveness.
+Analogous crisis-window panels for BTC-USD and Gold appear in Appendix B.
 6.1 Main Findings: Adaptiveness and Calibration
-The central finding of our study is twofold.
-First, static machine learning models like Quantile Regression, while producing sharp intervals, are fundamentally unreliable for risk forecasting due to their poor calibration.
-As shown in Table 1, the QR model consistently overshoots the 95% target, achieving  97% coverage across all assets.
-In risk management, such systematic miscalibration is a critical flaw, as it leads to an inefficient allocation of capital.
-Second, our proposed TCP framework demonstrates a clear and superior ability to adapt to changing market conditions.
-This is powerfully illustrated in Figure 1.
-During the low-volatility period of February 2020 for S & P 500, TCP’s interval is appropriately narrow.
-As the market crashes in March, the interval intelligently and dramatically expands to account for the heightened risk.
-Crucially, it begins to contract again in April as the market stabilizes.
-This dynamic behavior is precisely what is required of a robust risk management system.
-In contrast, the QR model’s interval remains static and inefficiently wide, while the GARCH and Historical Simulation models exhibit clumsy and less responsive adaptations.
-Similar visualizations for BTC-USD and Gold, which show consistent findings, are provided in Appendix B.
-While TCP’s empirical coverage of 86-88% undershoots the 95% target in this experiment, its adaptive mechanism, driven by the modified Robbins-Monro scheme, represents a principled attempt at calibration that is entirely absent in the static QR model.
-This ability to navigate the coverage-sharpness tradeoff in real-time is the core contribution of our framework.
-In summary, the results reveal that TCP is a more reliable and trustworthy tool for real-world financial forecasting, where adaptability to non-stationarity and adherence to specified risk tolerances are paramount.
-6.2 Hyperparameter Sensitivity
-To evaluate the robustness of TCP, we performed a sensitivity analysis on its key hyperparameters: the window size $w$ and the initial learning rate $\gamma_{0}$.
-We tested TCP on the S&P 500 dataset with window sizes $w\in\{100,252,500\}$ and learning rates $\gamma_{0}\in\{0.005,0.01,0.05\}$.
-The results are summarized in Table 2.
-This analysis was repeated for BTC-USD and Gold, yielding similar conclusions about the model’s stability; these additional results are provided in Appendix C.
-Table 2: TCP Sensitivity Analysis on S&P 500 Data
+Three findings emerge clearly from Table 1 and the crisis-window visualization (Fig. 1):
+  1. 1.
+    (i) Calibration vs.
+sharpness trade-off.
+The rolling ML baseline (QR) produces the narrowest intervals across all assets but is materially under-calibrated relative to the 95% target: SP500 86.3%, BTC 89.6%, Gold 87.6%.
+This under-coverage is the price for sharpness (e.g., SP500 average width 2.686 versus TCP’s 3.598).
+In other words, rolling QR concentrates mass too aggressively in volatile regimes.
+  2. 2.
+    (ii) TCP/TCP-RM achieve near-nominal coverage with substantially tighter bands than Historical Simulation.
+TCP coverage is close to the target on all assets (SP500 94.4%, BTC 95.9%, Gold 94.3%), while keeping intervals 30–38% narrower than Historical Simulation: SP500: 3.598 vs 5.058 (-28.9%); BTC: 11.235 vs 18.058 (-37.8%); Gold: 2.663 vs 4.024 (-33.8%).
+Thus, TCP delivers a better calibration–sharpness balance than Hist, which attains decent coverage but with overly conservative widths.
+  3. 3.
+    (iii) Robbins–Monro improves calibration at essentially no cost in width.
+TCP-RM yields a small positive lift in coverage relative to TCP with negligible width change on all assets (e.g., SP500: +0.28 percentage points coverage, +0.009 width).
+This supports the utility of the online offset for fine-tuning calibration without sacrificing sharpness.
+GARCH under-covers consistently ($\approx$0.83–0.85) while not being as sharp as QR, suggesting its conditional-Gaussian volatility path is too slow to adapt to return asymmetries and tail thickness during stress.
+Crisis-window behavior.
+Figure 1 (SP500, Feb–Apr 2020) shows how each method behaves during the shock.
+The red dots mark days where the realized return falls outside the reported 95% prediction interval (i.e., miscoverage events).
+TCP bands widen abruptly in early March, remain elevated through the peak volatility in mid–late March, and begin to contract in April as conditions stabilize.
+TCP and TCP-RM deliver similar behavior and, in this short 55-day window, both achieve coverage of $\approx 0.80$ with average widths $\approx 11.55$–11.57.
+Rolling QR widens only modestly, keeps much tighter bands (width =7.69=7.69), and shows the most red dots around the extremes, consistent with its lower window coverage (0.709).
+GARCH ramps up more smoothly, reflecting its volatility recursion, and lands at the same window coverage as TCP (0.800) with narrower bands (width =9.66=9.66).
+Historical Simulation reacts more slowly and is the most under-covered here (0.655), despite having the narrowest bands in this window (width =5.91=5.91); this lag is expected from its long empirical lookback.
+Overall, the crisis snapshot illustrates a general pattern under abrupt regime change: all methods temporarily under-cover, but TCP/TCP-RM adapt fastest by inflating bands right when volatility jumps, while QR remains sharp at the cost of additional misses.
+This visual evidence complements the full out-of-sample results in Table 1, where TCP/TCP-RM recover to near-nominal calibration over the full evaluation horizon.
+Analogous crisis-window panels for BTC-USD and Gold are reported in the Appendix B.
+Implementation note (comparability).
+All learning-based methods use the same feature set.
+TCP/TCP-RM are evaluated out-of-sample with a rolling window of 252 points split into 192 training and 60 calibration points.
+Rolling QR is also evaluated out-of-sample using a 192-point trailing training window.
+The slight differences in the “Predictions” counts (e.g., 1509 for QR vs.
+1449 for TCP/TCP-RM) reflect different warm-up requirements (QR does not reserve a calibration slice).
+Aligning the evaluation horizon to the longest warm-up yields qualitatively identical conclusions: QR is sharp but under-calibrated; TCP/TCP-RM are near-nominal and materially sharper than Historical Simulation.
+6.2 Hyperparameter Sensitivity (SP500, TCP-RM)
+We stress–test the adaptive layer by varying the rolling window $w$ and the initial step size $\gamma_{0}$ for TCP-RM(Table 2).
+For comparability we fix the calibration slice at $|\mathcal{C}|=40$ and sweep $w\in\{100,252,500\}$ and $\gamma_{0}\in\{0.005,0.01,0.05\}$ on SP500.
+Results for BTC-USD and Gold exhibit the same pattern and are reported in Appendix D.
+Table 2: TCP-RM sensitivity on S&P 500 with $|\mathcal{C}|=40$ (target 95%).
 Hyperparameters Performance
-Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
-100 0.005 0.8463 2.6521
-100 0.010 0.8569 2.6288
-100 0.050 0.8879 2.5847
-252 0.005 0.8496 2.7095
-252 0.010 0.8613 2.6611
-252 0.050 0.8958 2.5923
-500 0.005 0.8524 2.7712
-500 0.010 0.8631 2.7309
-500 0.050 0.8965 2.6614
-The analysis shows that TCP’s performance is remarkably stable.
-While a larger window size ($w=500$) slightly increases interval width as it incorporates more historical data, the coverage remains consistent.
-Similarly, a higher learning rate ($\gamma_{0}=0.05$) allows for faster adaptation, pushing the coverage closer to 90% without causing instability.
-Across all settings, TCP maintains a strong balance between coverage and sharpness, confirming that its effectiveness is not an artifact of fine-tuning.
-This robustness is a critical feature for practical deployment, as it reduces the need for costly hyperparameter optimization.
+Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
+100 0.005 0.9557 4.5188
+100 0.010 0.9557 4.5202
+100 0.050 0.9550 4.5315
+252 0.005 0.9469 3.4605
+252 0.010 0.9469 3.4673
+252 0.050 0.9524 3.4878
+500 0.005 0.9559 3.2627
+500 0.010 0.9659 3.2815
+500 0.050 0.9759 3.3970
+What we learn.
+(i) Increasing $w$ consistently sharpens intervals: widths fall from $\approx 4.52$ (w=100w{=}100) to 3.263.263.403.40 (w=500w{=}500), while coverage stays at or above the target band.
+(ii) At a fixed $w$, larger $\gamma_{0}$ nudges coverage upward with a small width trade-off; e.g., at w=252w{=}252, $\gamma_{0}{=}0.050$ lifts coverage from $0.9469$ to $0.9524$ (+0.55pp) with a $\sim 0.03$ increase in width; at w=500w{=}500, very large $\gamma_{0}$ overshoots the target (0.9759).
+Takeaway.
+TCP-RM is robust to reasonable choices of $(w,\gamma_{0})$.
+For the main experiments, we retain w=252w{=}252 and $\gamma_{0}{=}0.01$, which stay near the nominal level while keeping intervals tight; if one prefers even sharper intervals, w=500w{=}500 with $\gamma_{0}\in[0.005,0.01]$ is a good choice without material over-coverage.
+We also perform sensitivity analysis on TCP. The results are given in Appendix E.
+6.3 Visualization for TCP vs. TCP-RM on S&P 500 (95% intervals)
+In this section, we present the TCP vs TCP-RM trace plot for S&P 500.
+Figure 2: TCP vs.
+TCP-RM on S&P 500 (95% intervals).
+Top: daily returns with two‐sided prediction bands for TCP (blue) and TCP-RM (purple).
+Middle: 30-day rolling empirical coverage; the dashed line marks the 95% target.
+Bottom: evolution of conformal thresholds: split‐conformal level $C_{\text{split}}$, online Robbins–Monro offset $C_{\text{RM}}$, and the effective total $C_{\text{eff}}=C_{\text{split}}+C_{\text{RM}}$.
+During the COVID-19 crash the bands widen sharply, the rolling coverage briefly departs from the 95% target and then re-enters the band, and $C_{\text{eff}}$ spikes before decaying as conditions stabilize.
+In Fig.
+2, TCP and TCP-RM behave similarly in tranquil periods, but during stress the effective threshold $C_{\text{eff}}$ (bottom panel) rises quickly, producing wider bands (top) and steering the 30-day rolling coverage back toward the 95% target (middle).
+The Robbins–Monro offset $C_{\text{RM}}$ remains near zero except around regime shifts, where it provides the extra adjustment that closes the gap to target coverage.
+This visual evidence complements Table 1: on S&P 500, TCP-RM attains $\approx$95% coverage with only a marginal width increase relative to TCP
 7.  CONCLUSION AND FUTURE WORK
-In this paper, we introduced Temporal Conformal Prediction (TCP), a flexible and adaptive framework for generating prediction intervals in financial time series.
-Our method combines the strengths of quantile regression with conformal calibration, and crucially adapts over time using a decaying learning rate to correct for miscalibration.
-Unlike traditional models, TCP does not assume stationarity or rely on rigid parametric forms, making it well-suited to handle the complex, noisy, and often non-i.i.d.
-nature of financial returns.
-Unlike existing adaptive conformal approaches (Gibbs et al., 2021), TCP explicitly leverages temporal exchangeability and a modified Robbins-Monro scheme to handle the unique non-stationarity of financial time series.
-This makes it particularly suited for regulatory applications where robustness to market regime shifts is critical.
-We benchmarked TCP against three widely used alternatives: standard Quantile Regression (QR), GARCH models, and Historical Simulation.
-Our results, across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), highlight a crucial finding: static machine learning models like QR, while producing sharp intervals, are fundamentally miscalibrated and systematically overshoot the nominal 95% coverage target.
-In contrast, TCP’s adaptive mechanism demonstrates a clear and superior ability to adjust its intervals in response to market volatility, offering a more principled approach to achieving the desired coverage.
-Furthermore, a detailed sensitivity analysis confirmed that TCP’s performance is robust across a range of hyperparameter settings, underscoring the stability of the framework.
-One of the key takeaways is that TCP offers a principled way to navigate the coverage–sharpness tradeoff, which is central to real-world forecasting applications.
-In practice, achieving properly calibrated intervals is paramount for interpretability, regulatory compliance, and capital efficiency.
-TCP’s distribution-free design aligns with Basel III/IV and SEC stress-testing requirements, which mandate robust risk assessments under varying market conditions.
-By providing valid intervals without parametric assumptions, TCP could streamline compliance workflows while improving capital allocation efficiency.
-By design, TCP fills a growing need for distribution-free, adaptive forecasting tools in financial modeling, a high-impact domain given increased scrutiny around risk assessments and regulatory robustness.
+This paper introduced Temporal Conformal Prediction (TCP), a simple and effective framework for calibrated prediction intervals in nonstationary time series.
+TCP combines a rolling, exchangeability‐aware split–conformal layer with a modern quantile forecaster, and the TCP–RM variant adds a single Robbins–Monro offset that adjusts the conformal threshold online using observed coverage errors.
+Across three asset classes, TCP/TCP–RM deliver a strong calibration–sharpness balance: coverage is close to the 95% target while intervals are materially narrower than Historical Simulation.
+In contrast, the rolling QR baseline attains the narrowest bands but undershoots the target coverage, especially in turbulent regimes.
+Crisis‐window visualizations around the COVID‐19 crash illustrate TCP’s intended adaptiveness: intervals widen rapidly as volatility spikes and contract as conditions stabilize, with red dots indicating miscoverage events.
+A targeted sensitivity analysis shows robustness to reasonable choices of rolling window and Robbins–Monro step size, and suggests default settings that are stable in practice.
+Limitations.
+Like any sequential method, TCP can temporarily under‐cover during abrupt regime shifts before its calibration catches up.
+Performance also inherits the quality of the base quantile forecaster and the choice of windowing; extremely heavy tails or structural breaks can stress any fixed feature set.
+Finally, while TCP–RM improves calibration at negligible width cost on average, very aggressive step sizes can overcorrect.
+Practical implications.
+By providing distribution‐free intervals that adapt in time, TCP aligns naturally with backtesting and governance needs in risk management (coverage targets, exception tracking, and scenario robustness).
+The framework is lightweight (single threshold) and model‐agnostic (plugs into any quantile forecaster), making deployment straightforward in production pipelines.
 7.1 Future Work
-Building on the foundation established in this work, we identify several avenues for future exploration.
-While our study focuses on univariate daily returns, several promising extensions remain.
-First, TCP can be generalized to multivariate or portfolio‐level forecasting by modeling joint or conditional quantiles This is where deep copula models (Aich et al., 2025) could be used to capture complex, non-linear dependencies between assets.
-Second, richer quantile estimators, such as neural networks or advanced tree‐based methods using additional market features, may further improve interval sharpness and adaptability.
-Third, it would be valuable to benchmark TCP against volatility models that leverage realized‐volatility inputs, for example true Realized-GARCH fitted on intraday realized variances or Realized-GARCH variants using rolling-window proxies.
-Fourth, a particularly promising avenue is applying TCP to the detection of financial bubbles; the behavior of the interval bounds themselves could serve as a novel, real-time indicator of market exuberance, while the model’s standard crash signal could be used to forecast the subsequent implosion, providing a new perspective on the work of Phillips et al.
-from 2018.
-Furthermore, the framework can be adapted from general risk monitoring to event-specific forecasting, such as predicting the likelihood of extreme monthly drawdowns (e.g., $>$10% market declines) by aggregating the model’s daily signals.
-Finally, deploying TCP in live trading simulations or real-time risk‐management systems will help assess its operational robustness, latency, and scalability.
-We leave these directions for future exploration.
-Finally, the framework’s principles are highly applicable beyond finance, wherever calibrated uncertainty quantification is needed for non-stationary processes.
-In marketing, for instance, TCP could reframe forecasting as a risk management problem, providing robust prediction intervals for demand forecasting and inventory management, or quantifying the uncertain returns from marketing mix models and budget allocations.
-In biostatistics, its ability to handle non-stationarity is ideal for real-time patient monitoring in ICUs, forecasting blood-glucose ranges for chronic disease management, and adapting personalized drug dosing.
-Further applications exist in renewable energy for forecasting power output to ensure grid stability and in sports analytics for dynamic, in-game strategic decision-making and injury risk assessment.
-Exploring these domains would underscore TCP’s potential as a foundational tool for adaptive forecasting.
+Several extensions follow naturally:
+  1. 1.
+    Multivariate and portfolio TCP. Generalize to joint/conditional quantiles for baskets of assets; couple TCP with copula or factor models to control portfolio‐level coverage.
+  2. 2.
+    Conditional and localized coverage. Target coverage conditional on covariates (e.g., volatility state) and explore adaptive nonconformity scores that emphasize tail risk or signed errors.
+  3. 3.
+    Richer forecasters and features.
+Swap the base learner for neural quantile models or boosted variants with realized‐volatility inputs (intraday RV), order‐flow, or macro features; assess benefits for sharpness at fixed coverage.
+  4. 4.
+    Change–point aware adaptation. Trigger temporary step–size boosts or threshold projections when a regime break is detected to shorten the transient under/over‐shoot in coverage.
+  5. 5.
+    Heavy‐tail robustness. Investigate EVT‐inspired nonconformity scores or asymmetric penalties to better stabilize performance under extreme shocks.
+  6. 6.
+    Decision integration. Map TCP intervals to VaR/ES, margin, and capital rules; evaluate utility/capital efficiency vs. conservative baselines in realistic backtests.
+  7. 7.
+    Beyond finance. TCP’s adaptiveness is broadly applicable wherever nonstationarity and calibrated uncertainty matter:
+    • Biostatistics/medicine: ICU vitals monitoring and glucose‐range forecasting with patient–specific adaptation; adaptive dosing intervals under drift.
+    • Epidemiology: Calibrated intervals for nowcasts/forecasts of incidence and $R_{t}$ that widen under reporting delays or testing-policy shifts, plus adaptive risk bands for hospitalization/ICU load as transmission regimes change.
+    • Genetics/Genomics: Uncertainty bands for polygenic risk scores and eQTL effect sizes that remain calibrated under population/covariate shift; interval forecasts for gene expression (bulk or single-cell) with batch effects and sequencing-depth drift; calibrated pathogenicity scores for rare variants.
+    • Sports analytics: Game-state–aware intervals for win probability, expected goals/points, and player performance that adapt to tempo and lineup changes; injury-risk ranges for load management with rapid recalibration after shocks (e.g., in-game injuries or schedule congestion).
+    • Renewable energy: Wind/solar generation intervals that expand/contract with weather regimes to aid grid balancing.
+    • Operations/marketing: Demand and lead‐time intervals for inventory control; calibrated lift ranges in marketing‐mix models under campaign drift.
+    • Transportation and climate: Travel‐time intervals under traffic regime shifts; local hazard/risk bands for extreme weather.
+In sum, TCP offers a compact, distribution‐free, and adaptive route to calibrated uncertainty quantification in the presence of temporal drift.
+Its empirical performance, visual adaptiveness in crisis windows, and robustness to sensible hyperparameters suggest it is a practical default for risk forecasting and a promising building block for broader sequential decision systems.
+8.  DATA AND CODE AVAILABILITY
+The financial returns dataset and all code used in this paper are publicly available at: https://github.com/agnivibes/temporal-conformal-prediction-tcp.
 Appendix A MATHEMATICAL PROOFS
+This section provides the proofs of Theorems 4.1 and 4.2.
 A.1 Proof of Finite-Sample Validity (Theorem 4.1)
@@
 Therefore, the probability of miscoverage is at most $\alpha$, which implies that the probability of correct coverage is at least $1-\alpha$. This completes the proof.
-A.2 Proof of Asymptotic Coverage (Theorem 4.2)
-Theorem Statement.
-Under the online update $C_{t+1}=C_{t}+\gamma_{t}e_{t}$, where $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$, with step-sizes $\gamma_{t}$ satisfying the Robbins-Monro conditions ($\sum\gamma_{t}=\infty,\sum\gamma_{t}^{2}<\infty$), and assuming ergodicity, then almost surely:
-$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\alpha.$
-Proof. This proof relies on the standard convergence results for stochastic approximation algorithms, as established by Robbins and Monro (1951) and further developed by Kushner and Yin (2003).
-Let $\mathcal{F}_{t-1}$ be the filtration representing all information available up to time $t-1$. The update rule for our conformal threshold, $C_{t}$, is:
-$C_{t+1}=C_{t}+\gamma_{t}(\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha).$
-This is a stochastic approximation algorithm designed to find the root of the function $f(C)=\mathbb{E}[\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha|C_{t}=C]$.
-The algorithm seeks the value $C^{*}$ such that $f(C^{*})=0$.
-(i) Martingale Decomposition.
-The update can be decomposed into a deterministic component and a stochastic noise component. Let the coverage error be $e_{t}=\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha$. We can write:
-$C_{t+1}-C_{t}=\gamma_{t}\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]+\gamma_{t}(e_{t}-\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]).$
-The term $\Delta_{t}=e_{t}-\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]$ is a martingale-difference sequence, as $\mathbb{E}[\Delta_{t}|\mathcal{F}_{t-1}]=0$.
-This term represents the zero-mean noise driving the process.
-The term $\gamma_{t}\mathbb{E}[e_{t}|\mathcal{F}_{t-1}]$ is the “signal” that pushes $C_{t}$ towards the desired equilibrium.
-(ii) Almost Sure Convergence.
-The Robbins-Monro theorem guarantees that $C_{t}$ converges to the unique root $C^{*}$ almost surely, provided certain conditions are met. Our setup satisfies these conditions:
-  1. 1.
-    The learning rate $\gamma_{t}$ satisfies $\sum_{t=1}^{\infty}\gamma_{t}=\infty$ and $\sum_{t=1}^{\infty}\gamma_{t}^{2}<\infty$, by construction.
-  2. 2.
-    The error term $e_{t}$ is bounded, since $e_{t}\in[-\alpha,1-\alpha]$.
-  3. 3.
-    Under standard assumptions of ergodicity and continuity of the underlying return distribution, there exists a unique threshold $C^{*}$ such that the expected coverage error is zero: $\mathbb{E}[\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha|C_{t}=C^{*}]=0$.
-Given these conditions, standard stochastic approximation theory guarantees that $C_{t}\to C^{*}$ almost surely.
-(iii) Convergence of the Coverage Fraction.
-As $C_{t}$ converges to the stable point $C^{*}$, the system reaches an equilibrium where the expected coverage error is zero.
-By Kronecker’s lemma (Hall and Hyde, 1980), a consequence of the convergence of the stochastic approximation is that the long-run average of the error term also converges to zero:
-$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}e_{t}=0\quad a.s.$
-Substituting the definition of $e_{t}$:
-$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}(\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha)=0.$
-By the linearity of limits, this directly implies that the long-run miscoverage fraction converges to $\alpha$:
-$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])=\alpha.$
-The long-run fraction of observations falling inside the interval is simply one minus the miscoverage fraction. Therefore,
-$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\in[\ell_{t},u_{t}])=1-\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])=1-\alpha.$
-This completes the proof.
-Appendix B Appendix: Supplementary Visualizations
+A.2 Proof of Asymptotic Calibration (Theorem 4.2)
+Theorem.
+Under (A1)–(A4) and the update $C_{t+1}=C_{t}+\gamma_{t}e_{t}$ with $e_{t}=\mathbf{1}\{r_{t}\notin[\ell_{t},u_{t}]\}-\alpha$, we have $C_{t}\to C^{\star}$ a.s.
+where $g(C^{\star})=1-\alpha$, and $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{r_{t}\in[\ell_{t},u_{t}]\}\to 1-\alpha$ a.s.
+Proof.
+Step 1 (Notation and stability).
+Let $\mathcal{F}_{t}$ be the filtration generated by the data up to time $t$.
+Define $g(C)=\Pr\!\big(r_{t}\in[\ell_{t},u_{t}]\mid C_{t}=C\big)$ and $\phi(C)=\mathbb{E}[e_{t}\mid C_{t}=C]=\alpha-g(C)$.
+By (A3), $\phi$ is continuous and strictly decreasing in a neighborhood of its unique root $C^{\star}$ with $\phi(C^{\star})=0$.
+Under (A1) the two-sided nonconformity scores are bounded, hence there exists $S<\infty$ such that valid thresholds lie in $[0,S]$.
+Without loss of generality we assume the recursion is kept in a compact interval (e.g., via projection onto $[0,S]$); this is standard in SA and does not change the produced intervals.
+Step 2 (SA decomposition). Decompose the coverage error into signal + noise:
+$e_{t}\;=\;\mathbb{E}[e_{t}\mid\mathcal{F}_{t-1}]\;+\;\Delta_{t}\;=\;\phi(C_{t})\;+\;\Delta_{t},$
+where $\Delta_{t}:=e_{t}-\mathbb{E}[e_{t}\mid\mathcal{F}_{t-1}]$ is a martingale-difference sequence with $\mathbb{E}[\Delta_{t}\mid\mathcal{F}_{t-1}]=0$ and uniformly bounded second moments by (A2).
+The recursion becomes
+$C_{t+1}\;=\;C_{t}\;+\;\gamma_{t}\{\phi(C_{t})+\Delta_{t}\}.$ (3)
+Step 3 (Convergence of $C_{t}$ to $C^{\star}$). Consider the Lyapunov function $V(c)=(c-C^{\star})^{2}$. From (3),
+$C_{t+1}-C^{\star}\;=\;(C_{t}-C^{\star})+\gamma_{t}\{\phi(C_{t})+\Delta_{t}\},$
+hence
+$V(C_{t+1})$ $=(C_{t}-C^{\star})^{2}\;+\;2\gamma_{t}(C_{t}-C^{\star})\{\phi(C_{t})+\Delta_{t}\}\;+\;\gamma_{t}^{2}\{\phi(C_{t})+\Delta_{t}\}^{2}.$
+Taking conditional expectation and using $\mathbb{E}[\Delta_{t}\mid\mathcal{F}_{t-1}]=0$,
+$\mathbb{E}[V(C_{t+1})\mid\mathcal{F}_{t-1}]\;=\;V(C_{t})\;+\;2\gamma_{t}(C_{t}-C^{\star})\phi(C_{t})\;+\;\gamma_{t}^{2}\,\mathbb{E}\big[\{\phi(C_{t})+\Delta_{t}\}^{2}\mid\mathcal{F}_{t-1}\big].$
+By (A1)–(A2) and compactness of the state space, there exists $K<\infty$ with $\mathbb{E}[(\phi(C_{t})+\Delta_{t})^{2}\mid\mathcal{F}_{t-1}]\leq K$ a.s.
+By (A3), $\phi$ is strictly decreasing near $C^{\star}$, hence there exists $\eta>0$ and $\rho>0$ such that $(C-C^{\star})\phi(C)\leq-\eta(C-C^{\star})^{2}$ whenever $|C-C^{\star}|\leq\rho$.
+Since iterates remain in a compact interval and step-sizes decrease, standard SA stability (together with projection if used) ensures that eventually $|C_{t}-C^{\star}|\leq\rho$ a.s.
+(alternatively, one may argue by contradiction using the drift term).
+Therefore, for all large $t$,
+$\mathbb{E}[V(C_{t+1})\mid\mathcal{F}_{t-1}]\;\leq\;V(C_{t})\;-\;2\eta\,\gamma_{t}(C_{t}-C^{\star})^{2}\;+\;K\,\gamma_{t}^{2}.$
+By the Robbins–Siegmund supermartingale convergence lemma, since $\sum_{t}\gamma_{t}^{2}<\infty$ and $V(\cdot)\geq 0$, we obtain that $V(C_{t})$ converges a.s.
+and $\sum_{t}\gamma_{t}(C_{t}-C^{\star})^{2}<\infty$ a.s.
+Because $\sum_{t}\gamma_{t}=\infty$ (A4), the latter implies $\liminf_{t}(C_{t}-C^{\star})^{2}=0$, and together with the monotone drift near $C^{\star}$ yields $C_{t}\to C^{\star}$ almost surely.
+Step 4 (Time-average calibration).
+Recall $e_{t}=\phi(C_{t})+\Delta_{t}$.
+Since $C_{t}\to C^{\star}$ and $\phi$ is continuous, $\phi(C_{t})\to 0$ a.s.; by Cesàro averaging, $\frac{1}{T}\sum_{t=1}^{T}\phi(C_{t})\to 0$ a.s.
+By (A2) and square-integrability of $\Delta_{t}$, the strong law for martingale differences (e.g., Hall & Heyde, 1980) implies $\frac{1}{T}\sum_{t=1}^{T}\Delta_{t}\to 0$ a.s.
+Hence
+$\frac{1}{T}\sum_{t=1}^{T}e_{t}\;=\;\frac{1}{T}\sum_{t=1}^{T}\phi(C_{t})\;+\;\frac{1}{T}\sum_{t=1}^{T}\Delta_{t}\;\xrightarrow{\text{a.s.}}\;0.$
+Equivalently, $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{r_{t}\in[\ell_{t},u_{t}]\}\to 1-\alpha$ a.s., which completes the proof.
+Appendix B SUPPLEMENTARY VISUALIZATIONS
 This section provides the prediction interval visualizations for BTC-USD (3) and Gold (4) during the COVID-19 market crash (February-April 2020).
@@
 Figure 4: A comparison of 95% prediction intervals from all four models for Gold daily returns during the COVID-19 market crash (February-April 2020).
-Appendix C Appendix: Supplementary Sensitivity Analysis
-This section provides the full sensitivity analysis results for BTC-USD (3 and Gold (4), complementing the main analysis on the S&P 500 presented in the results section.
-The findings confirm that the TCP framework is robust and performs consistently across different asset classes and hyperparameter settings.
-Table 3: TCP Sensitivity Analysis on BTC-USD Data
+Appendix C VISUALIZATIONS OF THE TCP VS TCP-RM TRACE PLOTS
+This section provides the visualizations of the TCP vs TCP-RM trace plots for BTC-USD (Figure 5) and Gold (Figure 6), complementing the main visualization for the S&P 500 presented in the results section.
+Figure 5: TCP vs.
+TCP-RM on BTC-USD (95% intervals).
+Same layout as Fig.
+2.
+Large, frequent shocks lead to sustained but controlled elevations in $C_{\text{eff}}$, keeping rolling coverage close to 95% while allowing intervals to shrink as volatility recedes.
+Figure 6: TCP vs.
+TCP-RM on Gold (95% intervals).
+Same layout as Fig.
+2.
+Spikes in $C_{\text{eff}}$ are concentrated around crisis windows; elsewhere, $C_{\text{RM}}$ stays near zero and the method behaves like standard split-conformal TCP.
+Appendix D SUPPLEMENTARY SENSITIVITY ANALYSIS
+This section provides the full sensitivity analysis results for BTC-USD (Table 3) and Gold (Table 4), complementing the main analysis on the S&P 500 presented in the results section.
+The findings confirm that the TCP-RM framework is robust and performs consistently across different asset classes and hyperparameter settings.
+Table 3: TCP-RM sensitivity on BTC-USD with $|\mathcal{C}|=40$.
 Hyperparameters Performance
-Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
-100 0.005 0.8716 9.6583
-100 0.010 0.8806 9.6309
-100 0.050 0.9089 9.5891
-252 0.005 0.8751 9.6644
-252 0.010 0.8854 9.6109
-252 0.050 0.9165 9.5226
-500 0.005 0.8767 9.7543
-500 0.010 0.8861 9.7121
-500 0.050 0.9157 9.6235
-Table 4: TCP Sensitivity Analysis on Gold Data
+Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
+100 0.005 0.9519 17.1883
+100 0.010 0.9519 17.1899
+100 0.050 0.9519 17.2031
+252 0.005 0.9593 10.7970
+252 0.010 0.9593 10.7974
+252 0.050 0.9593 10.8004
+500 0.005 0.9517 9.0519
+500 0.010 0.9525 9.0544
+500 0.050 0.9542 9.0737
+Table 4: TCP-RM sensitivity on Gold with $|\mathcal{C}|=40$.
 Hyperparameters Performance
-Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
-100 0.005 0.8654 2.2101
-100 0.010 0.8751 2.1987
-100 0.050 0.9034 2.1643
-252 0.005 0.8703 2.2214
-252 0.010 0.8806 2.2021
-252 0.050 0.9096 2.1589
-500 0.005 0.8732 2.2456
-500 0.010 0.8821 2.2298
-500 0.050 0.9103 2.1887
+Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
+100 0.005 0.9538 3.7060
+100 0.010 0.9532 3.7066
+100 0.050 0.9532 3.7109
+252 0.005 0.9517 2.6241
+252 0.010 0.9524 2.6245
+252 0.050 0.9524 2.6294
+500 0.005 0.9534 2.4083
+500 0.010 0.9609 2.4212
+500 0.050 0.9542 2.4105
+Appendix E SUPPLEMENTARY SENSITIVITY ANALYSIS ON TCP (NO RM)
+For completeness, in Tables 5, 6 and 7, we also report sensitivity for TCP (split-conformal only, no Robbins–Monro offset). We vary $w\in\{100,252,500\}$ with $|\mathcal{C}|=40$ for each asset.
+Table 5: TCP sensitivity on S&P 500 with $|\mathcal{C}|=40$.
+Window Size ($w$) Coverage Interval Width
+100 0.9557 4.5174
+252 0.9455 3.4538
+500 0.9450 3.2376
+Table 6: TCP sensitivity on BTC-USD with $|\mathcal{C}|=40$.
+Window Size ($w$) Coverage Interval Width
+100 0.9519 17.1866
+252 0.9593 10.7967
+500 0.9500 9.0469
+Table 7: TCP sensitivity on Gold with $|\mathcal{C}|=40$.
+Window Size ($w$) Coverage Interval Width
+100 0.9538 3.7055
+252 0.9517 2.6234
+500 0.9417 2.3872
+As with TCP-RM, larger $w$ generally yields narrower intervals.
+Without the adaptive offset, coverage can dip slightly below $95\%$ in some settings, underscoring the benefit of the online calibration layer.

### v4

@@
 Three results are consistent across assets.
-First, rolling QR yields the sharpest intervals but is materially under‐calibrated (e.g., S&P 500: 86.3% vs.
+First, rolling QR yields the sharpest intervals but is materially under-calibrated (e.g., S&P 500: 83.2% vs.
 95% target).
-Second, TCP (and TCP–RM) achieves near‐nominal coverage while being substantially narrower than Historical Simulation (e.g., S&P 500: $29\%$ width).
-Third, the RM update improves calibration with negligible width cost.
+Second, TCP (and TCP–RM) achieves near-nominal coverage across assets, with intervals that are wider than Historical Simulation in this evaluation (e.g., S&P 500: 5.21 vs.
+5.06).
+Third, the RM update changes calibration and width only marginally at our default hyperparameters.
 Crisis‐window visualizations around March 2020 show TCP/TCP–RM expanding and then contracting their interval bands promptly as volatility spikes and recedes, with red dots marking days where realized returns fall outside the reported 95% interval (miscoverage).
@@
 Asset Model Empirical Coverage Avg. Interval Width Predictions
-SP500 TCP 0.9441 3.5975 1449
-TCP-RM 0.9469 3.6068 1449
-QR 0.8628 2.686 1509
+SP500 TCP 0.9523 5.2097 1448
+TCP-RM 0.9530 5.2171 1448
+QR 0.8322 3.3106 1508
 GARCH 0.8269 3.0505 1670
 Hist 0.9312 5.0575 1468
-BTC-USD TCP 0.9586 11.2346 1449
-TCP-RM 0.9593 11.2374 1449
-QR 0.8960 9.8446 1509
+BTC-USD TCP 0.9537 20.8879 1448
+TCP-RM 0.9537 20.8859 1448
+QR 0.8554 12.6764 1508
 GARCH 0.8533 11.3908 1670
 Hist 0.9441 18.0576 1468
-Gold TCP 0.9427 2.6630 1449
-TCP-RM 0.9448 2.6696 1449
-QR 0.8761 2.2394 1509
+Gold TCP 0.9427 4.7368 1448
+TCP-RM 0.9427 4.7438 1448
+QR 0.8302 2.8249 1508
 GARCH 0.8365 2.6185 1670
@@
 Analogous crisis-window panels for BTC-USD and Gold appear in Appendix B.
-6.1 Main Findings: Adaptiveness and Calibration
+Main Findings: Adaptiveness and Calibration
 Three findings emerge clearly from Table 1 and the crisis-window visualization (Fig. 1):
-  1. 1.
 (i) Calibration vs.
 sharpness trade-off.
-The rolling ML baseline (QR) produces the narrowest intervals across all assets but is materially under-calibrated relative to the 95% target: SP500 86.3%, BTC 89.6%, Gold 87.6%.
-This under-coverage is the price for sharpness (e.g., SP500 average width 2.686 versus TCP’s 3.598).
-In other words, rolling QR concentrates mass too aggressively in volatile regimes.
-  2. 2.
-    (ii) TCP/TCP-RM achieve near-nominal coverage with substantially tighter bands than Historical Simulation.
-TCP coverage is close to the target on all assets (SP500 94.4%, BTC 95.9%, Gold 94.3%), while keeping intervals 30–38% narrower than Historical Simulation: SP500: 3.598 vs 5.058 (-28.9%); BTC: 11.235 vs 18.058 (-37.8%); Gold: 2.663 vs 4.024 (-33.8%).
-Thus, TCP delivers a better calibration–sharpness balance than Hist, which attains decent coverage but with overly conservative widths.
-  3. 3.
-    (iii) Robbins–Monro improves calibration at essentially no cost in width.
-TCP-RM yields a small positive lift in coverage relative to TCP with negligible width change on all assets (e.g., SP500: +0.28 percentage points coverage, +0.009 width).
-This supports the utility of the online offset for fine-tuning calibration without sacrificing sharpness.
-GARCH under-covers consistently ($\approx$0.83–0.85) while not being as sharp as QR, suggesting its conditional-Gaussian volatility path is too slow to adapt to return asymmetries and tail thickness during stress.
+The rolling ML baseline (QR) yields the narrowest intervals across all assets but is materially under-calibrated relative to the 95% target: SP500 83.2%, BTC 85.5%, Gold 83.0%.
+This under-coverage is the price for sharpness (e.g., SP500 average width 3.311 versus TCP’s 5.210).
+(ii) TCP/TCP–RM achieve near-nominal coverage with wider bands than Historical Simulation.
+TCP coverage is close to the target on all assets (SP500 95.2%, BTC 95.4%, Gold 94.3%).
+Relative to Historical Simulation, the intervals are wider in this setting (SP500: 5.210 vs.
+5.058; BTC: 20.888 vs.
+18.058; Gold: 4.737 vs.
+4.024), reflecting stricter calibration under elevated volatility.
+(iii) Robbins–Monro adds negligible change at these hyperparameters.
+TCP–RM’s coverage and width are essentially identical to TCP across assets (coverage deltas $\leq$ 0.0007; width deltas $\leq$ 0.002), indicating the online offset has minimal incremental effect here.
+GARCH under-covers consistently ($\approx$0.83–0.85) while being sharper than the calibrated methods, consistent with its conditional-Gaussian specification being slow to adapt to asymmetric tails during stress.
 Crisis-window behavior.
 Figure 1 (SP500, Feb–Apr 2020) shows how each method behaves during the shock.
-The red dots mark days where the realized return falls outside the reported 95% prediction interval (i.e., miscoverage events).
-TCP bands widen abruptly in early March, remain elevated through the peak volatility in mid–late March, and begin to contract in April as conditions stabilize.
-TCP and TCP-RM deliver similar behavior and, in this short 55-day window, both achieve coverage of $\approx 0.80$ with average widths $\approx 11.55$–11.57.
-Rolling QR widens only modestly, keeps much tighter bands (width =7.69=7.69), and shows the most red dots around the extremes, consistent with its lower window coverage (0.709).
-GARCH ramps up more smoothly, reflecting its volatility recursion, and lands at the same window coverage as TCP (0.800) with narrower bands (width =9.66=9.66).
-Historical Simulation reacts more slowly and is the most under-covered here (0.655), despite having the narrowest bands in this window (width =5.91=5.91); this lag is expected from its long empirical lookback.
-Overall, the crisis snapshot illustrates a general pattern under abrupt regime change: all methods temporarily under-cover, but TCP/TCP-RM adapt fastest by inflating bands right when volatility jumps, while QR remains sharp at the cost of additional misses.
-This visual evidence complements the full out-of-sample results in Table 1, where TCP/TCP-RM recover to near-nominal calibration over the full evaluation horizon.
-Analogous crisis-window panels for BTC-USD and Gold are reported in the Appendix B.
+The red dots mark days where the realized return falls outside the reported 95% prediction interval (miscoverage).
+TCP/TCP–RM bands widen abruptly into the March spike and contract as conditions stabilize, whereas QR remains comparatively tight and under-covers around extremes.
+GARCH adapts smoothly via its volatility recursion, and Historical Simulation reacts with a lag due to its long lookback.
+Overall, all methods temporarily under-cover at the break, but TCP/TCP–RM inflate bands right when volatility jumps and recover calibration thereafter.
+Analogous crisis-window panels for BTC-USD and Gold are reported in Appendix B.
 Implementation note (comparability).
@@
 Rolling QR is also evaluated out-of-sample using a 192-point trailing training window.
-The slight differences in the “Predictions” counts (e.g., 1509 for QR vs.
+The slight differences in the “Predictions” counts (e.g., 1508 for QR vs.
 1448 for TCP/TCP-RM) reflect different warm-up requirements (QR does not reserve a calibration slice).
-Aligning the evaluation horizon to the longest warm-up yields qualitatively identical conclusions: QR is sharp but under-calibrated; TCP/TCP-RM are near-nominal and materially sharper than Historical Simulation.
+Aligning the evaluation horizon to the longest warm-up yields qualitatively identical conclusions: QR is sharp but under-calibrated; TCP/TCP-RM are near-nominal and, in this evaluation, wider than Historical Simulation.
 6.1 Hyperparameter Sensitivity (SP500, TCP-RM)
@@
 Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
-100 0.005 0.9557 4.5188
-100 0.010 0.9557 4.5202
-100 0.050 0.9550 4.5315
-252 0.005 0.9469 3.4605
-252 0.010 0.9469 3.4673
-252 0.050 0.9524 3.4878
-500 0.005 0.9559 3.2627
-500 0.010 0.9659 3.2815
-500 0.050 0.9759 3.3970
+100 0.005 0.9575 5.3052
+100 0.010 0.9575 5.3066
+100 0.050 0.9581 5.3172
+252 0.005 0.9579 5.1168
+252 0.010 0.9579 5.1191
+252 0.050 0.9586 5.1314
+500 0.005 0.9525 5.4135
+500 0.010 0.9525 5.4319
+500 0.050 0.9567 5.5348
 What we learn.
-(i) Increasing $w$ consistently sharpens intervals: widths fall from $\approx 4.52$ (w=100w{=}100) to 3.263.263.403.40 (w=500w{=}500), while coverage stays at or above the target band.
-(ii) At a fixed $w$, larger $\gamma_{0}$ nudges coverage upward with a small width trade-off; e.g., at w=252w{=}252, $\gamma_{0}{=}0.050$ lifts coverage from $0.9469$ to $0.9524$ (+0.55pp) with a $\sim 0.03$ increase in width; at w=500w{=}500, very large $\gamma_{0}$ overshoots the target (0.9759).
+(i) Across these settings, w=252w{=}252 yields the narrowest intervals (width $\approx 5.12$), while w=100w{=}100 and w=500w{=}500 are wider ($\approx 5.31$ and 5.415.415.535.53, respectively), with coverage remaining near or just above the 95% target for $w\in\{100,252\}$ and slightly lower for some w=500w{=}500 configurations.
+(ii) At a fixed $w$, increasing $\gamma_{0}$ produces only minor changes: coverage nudges up by at most a few tenths of a percentage point with a negligible width increase (e.g., at w=252w{=}252, $\gamma_{0}{=}0.050$ lifts coverage from $0.9579$ to $0.9586$ with a $\sim 0.012$ increase in width; at w=500w{=}500, $\gamma_{0}{=}0.050$ raises coverage from $0.9525$ to $0.9567$ while widening intervals modestly).
 Takeaway.
-TCP-RM is robust to reasonable choices of $(w,\gamma_{0})$.
-For the main experiments, we retain w=252w{=}252 and $\gamma_{0}{=}0.01$, which stay near the nominal level while keeping intervals tight; if one prefers even sharper intervals, w=500w{=}500 with $\gamma_{0}\in[0.005,0.01]$ is a good choice without material over-coverage.
+TCP–RM is robust over $(w,\gamma_{0})$ in this range.
+For the main experiments, w=252w{=}252 with $\gamma_{0}{=}0.01$ offers a good balance (coverage $\approx 0.9579$, width $\approx 5.1191$).
+If one prefers higher coverage at w=500w{=}500, a slightly larger $\gamma_{0}$ (e.g., $0.05$) helps, at the cost of modestly wider bands.
 We also perform sensitivity analysis on TCP. The results are given in Appendix E.
@@
 The Robbins–Monro offset $C_{\text{RM}}$ remains near zero except around regime shifts, where it provides the extra adjustment that closes the gap to target coverage.
-This visual evidence complements Table 1: on S&P 500, TCP-RM attains $\approx$95% coverage with only a marginal width increase relative to TCP
+This visual evidence complements Table 1: on S&P 500, TCP-RM attains $\approx$95% coverage with only a marginal width increase relative to TCP.
 7.  CONCLUSION AND FUTURE WORK
@@
 TCP combines a rolling, exchangeability‐aware split–conformal layer with a modern quantile forecaster, and the TCP–RM variant adds a single Robbins–Monro offset that adjusts the conformal threshold online using observed coverage errors.
-Across three asset classes, TCP/TCP–RM deliver a strong calibration–sharpness balance: coverage is close to the 95% target while intervals are materially narrower than Historical Simulation.
+Across three asset classes, TCP/TCP–RM delivers a strong calibration–sharpness balance: coverage is close to the 95% target; in our evaluation, the intervals are wider than Historical Simulation, reflecting stricter calibration during volatile periods.
 In contrast, the rolling QR baseline attains the narrowest bands but undershoots the target coverage, especially in turbulent regimes.
@@
 Its empirical performance, visual adaptiveness in crisis windows, and robustness to sensible hyperparameters suggest it is a practical default for risk forecasting and a promising building block for broader sequential decision systems.
-8.  DATA AND CODE AVAILABILITY
-The financial returns dataset and all code used in this paper are publicly available at: https://github.com/agnivibes/temporal-conformal-prediction-tcp.
 Appendix A MATHEMATICAL PROOFS
@@
 This section provides the prediction interval visualizations for BTC-USD (3) and Gold (4) during the COVID-19 market crash (February-April 2020).
-Figure 3: A comparison of 95% prediction intervals from all four models for BTC-USD daily returns during the COVID-19 market crash (February-April 2020).
-Figure 4: A comparison of 95% prediction intervals from all four models for Gold daily returns during the COVID-19 market crash (February-April 2020).
+Figure 3: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for BTC-USD daily returns during the COVID-19 market crash (February-April 2020).
+Figure 4: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for Gold daily returns during the COVID-19 market crash (February-April 2020).
 Appendix C VISUALIZATIONS OF THE TCP VS TCP-RM TRACE PLOTS
@@
 The findings confirm that the TCP-RM framework is robust and performs consistently across different asset classes and hyperparameter settings.
+For BTC-USD, widths decrease as $w$ grows (tightest at w=500w{=}500) while coverage remains near 95%–96%; $\gamma_{0}$ has negligible impact within a fixed $w$.
+For Gold, widths are smallest at w=100w{=}100 and increase with $w$; coverage hovers around the nominal level; $\gamma_{0}$ effects are minor at fixed $w$.
 Table 3: TCP-RM sensitivity on BTC-USD with $|\mathcal{C}|=40$.
@@
 Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
-100 0.005 0.9519 17.1883
-100 0.010 0.9519 17.1899
-100 0.050 0.9519 17.2031
-252 0.005 0.9593 10.7970
-252 0.010 0.9593 10.7974
-252 0.050 0.9593 10.8004
-500 0.005 0.9517 9.0519
-500 0.010 0.9525 9.0544
-500 0.050 0.9542 9.0737
+100 0.005 0.9575 21.7806
+100 0.010 0.9575 21.7819
+100 0.050 0.9575 21.7926
+252 0.005 0.9572 21.6138
+252 0.010 0.9572 21.6154
+252 0.050 0.9572 21.6285
+500 0.005 0.9608 20.5050
+500 0.010 0.9608 20.5053
+500 0.050 0.9608 20.5071
 Table 4: TCP-RM sensitivity on Gold with $|\mathcal{C}|=40$.
@@
 Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
-100 0.005 0.9538 3.7060
-100 0.010 0.9532 3.7066
-100 0.050 0.9532 3.7109
-252 0.005 0.9517 2.6241
-252 0.010 0.9524 2.6245
-252 0.050 0.9524 2.6294
-500 0.005 0.9534 2.4083
-500 0.010 0.9609 2.4212
-500 0.050 0.9542 2.4105
+100 0.005 0.9494 4.6437
+100 0.010 0.9494 4.6460
+100 0.050 0.9494 4.6643
+252 0.005 0.9461 4.7812
+252 0.010 0.9461 4.7868
+252 0.050 0.9468 4.8285
+500 0.005 0.9500 5.1058
+500 0.010 0.9500 5.1188
+500 0.050 0.9542 5.2031
 Appendix E SUPPLEMENTARY SENSITIVITY ANALYSIS ON TCP (NO RM)
@@
 Window Size ($w$) Coverage Interval Width
-100 0.9557 4.5174
-252 0.9455 3.4538
-500 0.9450 3.2376
+100 0.9575 5.3038
+252 0.9572 5.1145
+500 0.9508 5.3951
 Table 6: TCP sensitivity on BTC-USD with $|\mathcal{C}|=40$.
 Window Size ($w$) Coverage Interval Width
-100 0.9519 17.1866
-252 0.9593 10.7967
-500 0.9500 9.0469
+100 0.9575 21.7792
+252 0.9572 21.6121
+500 0.9608 20.5048
 Table 7: TCP sensitivity on Gold with $|\mathcal{C}|=40$.
 Window Size ($w$) Coverage Interval Width
-100 0.9538 3.7055
-252 0.9517 2.6234
-500 0.9417 2.3872
-As with TCP-RM, larger $w$ generally yields narrower intervals.
-Without the adaptive offset, coverage can dip slightly below $95\%$ in some settings, underscoring the benefit of the online calibration layer.
+100 0.9487 4.6414
+252 0.9454 4.7756
+500 0.9492 5.0928
+As with TCP-RM, the effect of $w$ is asset-dependent (e.g., BTC narrows as $w$ increases; SP500 is tightest near w=252w{=}252; Gold widens with larger $w$).
+Without the adaptive offset, coverage can dip slightly below 95% in some settings, underscoring the benefit of the online calibration layer.

### v5

@@
 Abstract
-We propose Temporal Conformal Prediction (TCP), a distribution‐free framework for constructing well‐calibrated prediction intervals in nonstationary time series.
-TCP couples a modern quantile forecaster with a split‐conformal calibration layer on a rolling window and, in its TCP–RM variant, augments the conformal threshold with a single online Robbins–Monro (RM) offset to steer coverage toward a target level in real time.
-We benchmark TCP against GARCH, Historical Simulation, and a rolling Quantile Regression (QR) baseline across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold).
+We propose Temporal Conformal Prediction (TCP), a distribution-free framework for constructing well-calibrated prediction intervals in nonstationary time series.
+TCP couples a modern quantile forecaster with a split-conformal calibration layer on a rolling window and, in its TCP-RM variant, augments the conformal threshold with a single online Robbins-Monro (RM) offset to steer coverage toward a target level in real time.
+We benchmark TCP against GARCH, Historical Simulation, a rolling tree-based Quantile Regression (QR) model, a classical linear quantile regression baseline (QR-Linear), and an adaptive conformal method (ACI) across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold).
 Three results are consistent across assets.
-First, rolling QR yields the sharpest intervals but is materially under-calibrated (e.g., S&P 500: 83.2% vs.
-95% target).
-Second, TCP (and TCP–RM) achieves near-nominal coverage across assets, with intervals that are wider than Historical Simulation in this evaluation (e.g., S&P 500: 5.21 vs.
+First, both QR and QR-Linear yield the sharpest intervals but are materially under-calibrated, and even ACI remains below the nominal 95% target in our full-sample backtests.
+Second, TCP (and TCP-RM) achieves near-nominal coverage across assets, with intervals that are wider than Historical Simulation in this evaluation (e.g., S&P 500: 5.21 vs.
 5.06).
 Third, the RM update changes calibration and width only marginally at our default hyperparameters.
-Crisis‐window visualizations around March 2020 show TCP/TCP–RM expanding and then contracting their interval bands promptly as volatility spikes and recedes, with red dots marking days where realized returns fall outside the reported 95% interval (miscoverage).
-A sensitivity study confirms robustness to window size and step‐size choices.
+Crisis-window visualizations around March 2020 show TCP/TCP-RM expanding and then contracting their interval bands promptly as volatility spikes and recedes, with red dots marking days where realized returns fall outside the reported 95% interval (miscoverage).
+A sensitivity study confirms robustness to window size and step-size choices.
 Overall, TCP provides a practical, theoretically grounded solution to calibrated uncertainty quantification under distribution shift, bridging statistical inference and machine learning for risk forecasting.
@@
 1.  INTRODUCTION
-Financial risk estimation is more than a regulatory checkbox; it is foundational for market stability and investor confidence (Markowitz, 1952; J.P.
-Morgan/Reuters, 1996).
+Financial risk estimation is more than a regulatory checkbox; it is foundational for market stability and investor confidence (Markowitz1952; Morgan1996).
 Yet, when markets enter turbulent regimes, traditional risk models often fall short.
-Early implementations of Value-at-Risk (VaR) were predominantly parametric, often assuming (conditional) normality (Jorion, 2007; Dowd, 1998), assumptions that can systematically understate tail risk.
-Coherent alternatives such as Conditional Value-at-Risk (CVaR) were introduced to address some of these limitations (Rockafellar and Uryasev, 2000).
-Events like the 2008 global financial crisis (Brunnermeier, 2009) and the March 2020 stock-market crash (Mazur et al., 2021) exposed the fragility of parametric assumptions, especially in the tails.
+Early implementations of Value-at-Risk (VaR) were predominantly parametric, often assuming (conditional) normality (Jorion2007; Dowd1998), assumptions that can systematically understate tail risk.
+Coherent alternatives such as Conditional Value-at-Risk (CVaR) were introduced to address some of these limitations (Rockafellar2000).
+Events like the 2008 global financial crisis (Brunnermeier2009) and the March 2020 stock-market crash (Mazur2021) exposed the fragility of parametric assumptions, especially in the tails.
 The core challenge is that financial returns violate the independent and identically distributed (i.i.d.) assumption that underpins many statistical learning techniques.
-Real-world returns are heteroskedastic, exhibit regime shifts, and often have heavy-tailed distributions (Cont, 2001).
-GARCH-type models (Engle, 1982; Bollerslev, 1986) can capture some volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
+Real-world returns are heteroskedastic, exhibit regime shifts, and often exhibit heavy tails (Cont2001).
+GARCH-type models (Engle1982; Bollerslev1986) can capture some volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
 In recent years, conformal prediction has gained traction as a powerful distribution-free framework for quantifying uncertainty.
-Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk et al., 2005).
+Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk2005).
 However, these guarantees rest on exchangeability, a condition often violated by time series with temporal dependencies.
-Several works extend conformal methods to sequential or temporally dependent settings (Xu and Xie, 2021; Stankeviciute et al., 2021; Gibbs and Candès, 2021), but a robust and practical solution for financial markets remains an open challenge.
+Several works extend conformal methods to sequential or temporally dependent settings (Xu2021; Stankeviciute2021; Gibbs2021), but a robust and practical solution for financial markets remains an open challenge.
 This paper closes that gap.
@@
 Crucially, this architecture bridges statistical inference and machine learning, combining the theoretical rigor of conformal methods with the flexibility of a data-driven approach that can adapt to non-stationarity, volatility clustering, and abrupt market shifts.
-We benchmark TCP across three major asset classes, equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), against established models like GARCH, Historical Simulation, and standard Quantile Regression (QR).
-Our results highlight a critical flaw in static ML approaches: while QR produces sharp intervals, it is poorly calibrated and systematically misses the 95% target.
-By contrast, TCP demonstrates superior adaptive capabilities, adjusting its intervals in response to market volatility to better align with the desired coverage rate.
+We benchmark TCP across three major asset classes, equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), against established models like GARCH, Historical Simulation, a tree-based Quantile Regression (QR) forecaster, a classical linear quantile regression baseline (QR-Linear), and an adaptive conformal method (ACI).
+Our results highlight a critical flaw in static ML approaches: both QR and QR-Linear produce sharp intervals but are systematically under-calibrated, and ACI improves calibration only partially.
+By contrast, TCP demonstrates superior adaptive capabilities, adjusting its intervals in response to market volatility so that empirical coverage stays close to the 95% target.
 This adaptive property, which we visualize during the March 2020 crash, makes TCP a more reliable and principled framework for real-world risk management.
@@
 Traditional Financial Risk Models.
-Quantitative risk management traces its roots to portfolio theory (Markowitz, 1952) and was formalized around the VaR concept with J.P.
-Morgan’s RiskMetrics (J.P.
-Morgan/Reuters, 1996).
-In practice, three VaR paradigms have dominated: parametric approaches (often assuming conditional normality), non-parametric Historical Simulation, and Monte Carlo methods (Jorion, 2007; Dowd, 1998; Glasserman, 2004).
-Beyond VaR, CVaR provided a coherent alternative for tail-risk control (Rockafellar and Uryasev, 2000).
-However, many traditional approaches underperform in crisis backtests (Kupiec, 1995; Christoffersen, 1998), highlighting a blind spot when tail events dominate.
+Quantitative risk management traces its roots to portfolio theory (Markowitz1952) and was formalized around the VaR concept with J.P.
+Morgan’s RiskMetrics (Morgan1996).
+In practice, three VaR paradigms have dominated: parametric approaches (often assuming conditional normality), non-parametric Historical Simulation, and Monte Carlo methods (Jorion2007; Dowd1998; Glasserman2004).
+Beyond VaR, CVaR provided a coherent alternative for tail-risk control (Rockafellar2000).
+However, many traditional approaches underperform in crisis backtests (Kupiec1995; Christoffersen1998), highlighting a blind spot when tail events dominate.
 Advanced Time-Series and Machine-Learning Models.
-To better capture time-varying variance and volatility clustering, the GARCH family became foundational (Engle, 1982; Bollerslev, 1986); asymmetric variants such as EGARCH and GJR-GARCH address leverage and sign effects (Nelson, 1991; Glosten et al., 1993).
-For heavy-tailed markets, extreme value theory (EVT) offers principled tail modeling (III, 1975; McNeil and Frey, 2000), and VaR under heavy tails raises specific concerns (Danielsson and de Vries, 2000).
-On the machine-learning side, quantile regression directly targets conditional quantiles (Koenker and Bassett, 1978), with ensemble and neural variants such as quantile regression forests (Meinshausen, 2006) and quantile networks (Taylor, 2000); deep LSTMs capture complex temporal dependencies (Fischer and Krauss, 2018).
-We also emphasize the computational efficiency of gradient-boosted trees via LightGBM for quantile forecasting (Ke et al., 2017).
+To better capture time-varying variance and volatility clustering, the GARCH family became foundational (Engle1982; Bollerslev1986); asymmetric variants such as EGARCH and GJR-GARCH address leverage and sign effects (Nelson1991; Glosten1993).
+For heavy-tailed markets, extreme value theory (EVT) offers principled tail modeling (Pickands1975; McNeil2000), and VaR under heavy tails raises specific concerns (Danielsson2000).
+On the machine-learning side, quantile regression directly targets conditional quantiles (Koenker1978), with ensemble and neural variants such as quantile regression forests (Meinshausen2006) and quantile networks (Taylor2000); deep LSTMs capture complex temporal dependencies (Fischer2018).
+We also emphasize the computational efficiency of gradient-boosted trees via LightGBM for quantile forecasting (Ke2017).
 Conformal Prediction Theory.
-Conformal prediction provides distribution-free, finite-sample prediction sets under exchangeability (Vovk et al., 2005).
-Recent advances relax these assumptions for sequential or non-stationary settings through adaptive calibration and time-series‐aware procedures (Gibbs and Candès, 2021; Xu and Xie, 2021; Stankeviciute et al., 2021).
+Conformal prediction provides distribution-free, finite-sample prediction sets under exchangeability (Vovk2005).
+Recent advances relax these assumptions for sequential or non-stationary settings through adaptive calibration and time-series‐aware procedures (Gibbs2021; Xu2021; Stankeviciute2021).
 Our work builds on these foundations with a practical adaptive conformal framework tailored to financial risk forecasting, combining local split-conformal calibration with an online Robbins–Monro offset for temporal adaptation.
+Adaptive and Time-Series Conformal Prediction.
+Recent research has addressed non-stationary environments by adapting the conformal guarantee.
+Gibbs2021 and zaffran2022adaptive introduce Adaptive Conformal Inference (ACI), which updates the target miscoverage rate ($\alpha_{t}$) online based on past coverage errors.
+While effective, we argue that in heavy-tailed financial settings, oscillating the target probability can lead to unstable interval widths; TCP-RM instead adapts the conformal threshold ($C_{t}$) directly via a Robbins-Monro update, offering a stable, scalar lever for width control.
+Other approaches, such as CPTD (lin2022temporal), leverage cross-sectional exchangeability across a panel of time series to restore validity.
+In contrast, TCP is designed for the univariate setting (single asset) where cross-sectional pooling (e.g., mixing Bitcoin and Gold) is inappropriate due to distinct, idiosyncratic asset dynamics.
 3.  NOTATION
@@
 We consider two closely related procedures.
-TCP denotes split–conformal prediction on a rolling window: the window $\mathcal{W}_{t}$ of size $w$ is split into a training slice $\mathcal{T}_{t}$ (size $w_{\mathrm{tr}}$) and a calibration slice $\mathcal{C}_{t}$ (size $w_{\mathrm{cal}}$), $w_{\mathrm{tr}}+w_{\mathrm{cal}}=w$.
-Quantile forecasters $\widehat{q}_{\alpha/2},\widehat{q}_{1-\alpha/2}$ are fit on $\mathcal{T}_{t}$ only and the two-sided split–conformal threshold $C_{t}$ is computed from $\mathcal{C}_{t}$, yielding $\ell_{t+1},u_{t+1}]=[\widehat{q}_{\alpha/2}(X_{t+1})-C_{t},\ \widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}]$.
-TCP-RM augments TCP with an online Robbins–Monro (RM) offset $C_{t}^{\mathrm{RM}}$ updated by $C_{t+1}^{\mathrm{RM}}=C_{t}^{\mathrm{RM}}+\gamma_{t}\{\mathbf{1}(r_{t}\notin[\ell_{t},u_{t}])-\alpha\}$, with $\gamma_{t}=\gamma_{0}(1+\lambda t)^{-\beta}$.
-The effective threshold is $C_{t}^{\mathrm{eff}}=C_{t}+C_{t}^{\mathrm{RM}}$ and the interval becomes $\widehat{q}_{\alpha/2}(X_{t+1})-C_{t}^{\mathrm{eff}},\ \widehat{q}_{1-\alpha/2}(X_{t+1})+C_{t}^{\mathrm{eff}}]$.
-For comparability, the QR baseline is evaluated in the same walk-forward fashion using only the training slice ($|\mathcal{T}_{t}|=w_{\mathrm{tr}}$) and no calibration slice (rolling, out-of-sample).
+TCP applies split–conformal prediction on a rolling window, and TCP-RM augments TCP with a Robbins–Monro offset applied to the conformal threshold based on past coverage errors.
+Formal definitions and update equations are given in Section 4.4.
 4.  MATHEMATICAL BACKGROUND
 This section lays the theoretical groundwork for our proposed method.
+Our calibration mechanism is related in spirit to adaptive conformal procedures for nonstationary settings (zaffran2022adaptive; lin2022temporal) but differs in using a Robbins–Monro update on the conformal threshold while holding the nominal level fixed.
 We begin by formally defining the problem of prediction interval forecasting, then review the principles of classical conformal prediction and its limitations, and finally develop the core mathematical components of our Temporal Conformal Prediction (TCP) framework and its adaptive calibration mechanism.
@@
 Under (A1)–(A4), TCP provides local finite-sample validity on $\mathcal{C}_{t}$; TCP-RM adds asymptotic calibration in time via (1).
+Clarifications on Assumptions (A1–A4).
+To make the role of the assumptions explicit in the analysis and in Theorem 4.2, we record the following clarifications:
+  1. 1.
+    Stochastic approximation framework.
+The Robbins-Monro update for the offset is analyzed using classical stochastic-approximation arguments under mixing and bounded-variance noise; our assumptions instantiate those conditions in this setting ((Kushner2003)).
+  2. 2.
+    TCP-RM guarantees. Under (A1)–(A4), the time-average coverage converges to the target level and the effective threshold converges to a fixed point (Theorem 4.2).
+  3. 3.
+    TCP baseline (split-conformal).
+TCP uses split-conformal calibration within a rolling window; this is a local working approximation to exchangeability, and we do not claim global finite-sample validity for the full time series.
+  4. 4.
+    Assumption verification in practice.
+We empirically probe the mixing/independence assumptions using standard risk diagnostics (Kupiec1995 and Christoffersen1998 backtests); the lack of serial dependence in exceedances supports the validity of these conditions on our financial datasets.
 4.5 Proof of Asymptotic Calibration
@@
   • Lagged Returns: $\{r_{t-k}\}_{k=1}^{5}$ to capture short-term momentum and autoregressive effects.
-  • Rolling Volatility: $\sigma_{t}=\sqrt{\frac{1}{20-1}\sum_{k=1}^{20}\big(r_{t-k}-\bar{r}_{t,20}\big)^{2}}\quad\text{(rolling sample std, ddof=1)}$ where $\bar{r}_{t,20}=\frac{1}{20}\sum_{k=1}^{20}r_{t-k}$.
+  • Rolling Volatility:
+    $\sigma_{t}$ $=\sqrt{\frac{1}{20-1}\sum_{k=1}^{20}(r_{t-k}-\bar{r}_{t,20})^{2}},$ (3)
+    $\bar{r}_{t,20}$ $=\frac{1}{20}\sum_{k=1}^{20}r_{t-k}.$ (4)
+    Here $\sigma_{t}$ is the rolling sample standard deviation of the past 20 returns, computed with denominator 20120-1 (in pandas this corresponds to .rolling(20).std(ddof=1)).
   • Nonlinear Transformations: The squared prior return, $r_{t-1}^{2}$, and its sign, $\mathrm{sign}(r_{t-1})$, to account for non-linear dependencies and leverage effects.
-In implementation we use the rolling sample standard deviation via pandas (.rolling(20).std, ddof=1=1).
 5.2 Forecasting Models
@@
   2. 2.
+    Linear Quantile Regression (QR-Linear): Classical linear quantile regression trained with the pinball loss at $\alpha/2$ and $1-\alpha/2$, using the same features and the same rolling, out-of-sample protocol as QR (window $w_{\mathrm{tr}}{=}192$; no calibration slice).
+This baseline is included in the full-sample numerical benchmarks (Table 1), but we omit it from the crisis-window visualizations to avoid clutter; its behavior is intermediate between QR and ACI.
+  3. 3.
+    Adaptive Conformal Inference (ACI): A rolling implementation of Gibbs2021 that uses the same window split as TCP ($w_{\mathrm{tr}}=192$ training, $w_{\mathrm{cal}}=60$ calibration).
+However, instead of updating the conformal threshold $C_{t}$ (like TCP-RM), ACI updates the target miscoverage rate $\alpha_{t}$ online based on the most recent coverage error (step size $\eta=0.05$).
+  4. 4.
     Fixed-parameter GARCH(1,1) (EWMA): Conditional variance updated recursively
-    $\sigma_{t}^{2}=\omega+\alpha r_{t-1}^{2}+\beta\,\sigma_{t-1}^{2},\qquad(\omega,\alpha,\beta)=(10^{-6},\,0.05,\,0.9),$
+    $\sigma_{t}^{2}$ $=\omega+\alpha r_{t-1}^{2}+\beta\,\sigma_{t-1}^{2},$ (5)
+    $(\omega,\alpha,\beta)$ $=(10^{-6},\,0.05,\,0.9),$ (6)
     with coefficients held fixed and updated sequentially using only past data (no rolling re-estimation).
 One-step-ahead $(1-\alpha)$ intervals assume conditional normality with zero mean, i.e., $z_{1-\alpha/2}\sigma_{t},\;z_{1-\alpha/2}\sigma_{t}\,]$.
-  3. 3.
+  5. 5.
     Historical Simulation (Hist): Non-parametric empirical $(\alpha/2,\,1-\alpha/2)$ quantiles from a rolling 252-day window.
 5.3 Evaluation Framework
-To provide a rigorous comparison, we must define clear success criteria.
-This subsection outlines the key performance metrics, empirical coverage and interval sharpness, that we use to evaluate and contrast the models.
-We use two primary metrics:
+To provide a rigorous comparison, we define clear success criteria for all models.
+Our primary metrics are (i) marginal empirical coverage of the prediction intervals and (ii) average interval width (sharpness).
+In addition, we apply standard VaR backtests on the 5% lower tail implied by the 95% two-sided intervals.
+We use two core summary metrics:
   • Empirical Coverage: The proportion of out-of-sample observations that fall within their respective prediction intervals, $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{r_{t}\in[\ell_{t},u_{t}]\}$.
@@
 Sharper (narrower) intervals are preferred, provided that the target coverage is met.
-Our analysis focuses on the trade-off between these two metrics, as a model is only useful if it is both well-calibrated (correct coverage) and sharp (informative).
+VaR-style backtests.
+For each model and asset, we also perform standard 1-day-ahead VaR backtests on the 5% lower tail.
+Specifically, we report: (i) the unconditional coverage test of Kupiec1995 (LRuc), (ii) the independence test of Christoffersen1998 (LRind), and (iii) their sum $_{\mathrm{cc}}{=}LR$$_{\mathrm{uc}}{+}$LRind for conditional coverage.
+Each test yields a likelihood-ratio statistic and a $p$-value, together with the number of exceedances $x$ (days where the realized return falls below the nominal 5% VaR) out of $n$ out-of-sample predictions.
+Our empirical analysis therefore evaluates models along two dimensions, i.e., global calibration–sharpness, via empirical coverage and average width over the full sample, and risk-management diagnostics in the left tail, via the Kupiec/Christoffersen backtests.
 6.  RESULTS AND DISCUSSION
-We now assess Temporal Conformal Prediction (TCP) and our main contribution TCP-RM (TCP with a Robbins–Monro online calibration layer) against strong baselines (QR, GARCH, Historical Simulation) on three asset classes: S&P 500, BTC-USD, and Gold.
+We now assess Temporal Conformal Prediction (TCP) and our main contribution TCP-RM (TCP with a Robbins–Monro online calibration layer) against a set of baselines (ACI, QR, QR-Linear, GARCH, Historical Simulation) on three asset classes: S&P 500, BTC-USD, and Gold.
 The goal is to test whether the online calibration delivers near-nominal coverage under distribution shift while preserving sharpness.
@@
 We also report a hyperparameter sensitivity study (window size $w$, stepsize $\gamma_{0}$) to assess robustness of TCP-RM.
+QR-Linear appears in the full-sample benchmark table but is omitted from the crisis-window plots and window-level summary tables to keep the figures legible; numerically, its coverage and width lie between QR and ACI.
 Table 1: Model Performance Across Assets (Target Coverage: 95%)
@@
 TCP-RM 0.9530 5.2171 1448
+ACI 0.8736 3.7894 1448
 QR 0.8322 3.3106 1508
+QR-Linear 0.8846 4.2746 1508
 GARCH 0.8269 3.0505 1670
@@
 TCP-RM 0.9537 20.8859 1448
+ACI 0.8874 16.3529 1448
 QR 0.8554 12.6764 1508
+QR-Linear 0.8952 17.4916 1508
 GARCH 0.8533 11.3908 1670
@@
 TCP-RM 0.9427 4.7438 1448
+ACI 0.8294 2.9901 1448
 QR 0.8302 2.8249 1508
+QR-Linear 0.8932 3.9780 1508
 GARCH 0.8365 2.6185 1670
 Hist 0.9326 4.0238 1468
-Figure 1: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for S&P 500 daily returns during the COVID-19 market crash (Feb–Apr 2020).
+Figure 1: A comparison of 95% prediction intervals from six models (TCP, TCP-RM, QR, GARCH, Hist, ACI) for S&P 500 daily returns during the COVID-19 market crash (Feb–Apr 2020).
 Shaded bands show the interval; the red dots mark days where the realized return falls outside the 95% interval (miscoverage).
-TCP/TCP-RM bands widen rapidly into the March spike and contract in April, illustrating adaptiveness.
-Analogous crisis-window panels for BTC-USD and Gold appear in Appendix B.
+TCP/TCP-RM bands widen rapidly into the March spike and contract in April, while ACI maintains relatively narrow bands and under-covers more frequently.
+Analogous crisis-window panels for BTC-USD and Gold appear in Appendix E.
 Main Findings: Adaptiveness and Calibration
-Three findings emerge clearly from Table 1 and the crisis-window visualization (Fig. 1):
+Three findings emerge from Table 1 and the crisis-window visualization (Fig. 1):
 (i) Calibration vs.
-sharpness trade-off.
-The rolling ML baseline (QR) yields the narrowest intervals across all assets but is materially under-calibrated relative to the 95% target: SP500 83.2%, BTC 85.5%, Gold 83.0%.
-This under-coverage is the price for sharpness (e.g., SP500 average width 3.311 versus TCP’s 5.210).
+sharpness trade-off across learning baselines.
+The rolling ML baselines (QR, QR-Linear, ACI) yield substantially narrower intervals than TCP/TCP-RM, but remain under-calibrated relative to the 95% target, which is consistent with the VaR backtests where only TCP and TCP-RM are not rejected.
+Among these baselines, QR-Linear and ACI partially improve calibration over tree-based QR at the price of somewhat wider bands, yet all three still fall short of nominal coverage.
 (ii) TCP/TCP–RM achieve near-nominal coverage with wider bands than Historical Simulation.
-TCP coverage is close to the target on all assets (SP500 95.2%, BTC 95.4%, Gold 94.3%).
-Relative to Historical Simulation, the intervals are wider in this setting (SP500: 5.210 vs.
-5.058; BTC: 20.888 vs.
-18.058; Gold: 4.737 vs.
-4.024), reflecting stricter calibration under elevated volatility.
+TCP coverage is close to the 95% target on all assets (SP500 95.2%, BTC-USD 95.4%, Gold 94.3%).
+Relative to Historical Simulation, TCP/TCP-RM intervals are systematically wider (SP500: 5.21 vs.
+5.06; BTC-USD: 20.89 vs.
+18.06; Gold: 4.74 vs.
+4.02), reflecting stricter calibration under elevated volatility.
+Historical Simulation attains high but slightly sub-nominal coverage (93–94%) with narrower bands, consistent with its reliance on a long empirical window that smooths abrupt regime changes.
 (iii) Robbins–Monro adds negligible change at these hyperparameters.
-TCP–RM’s coverage and width are essentially identical to TCP across assets (coverage deltas $\leq$ 0.0007; width deltas $\leq$ 0.002), indicating the online offset has minimal incremental effect here.
-GARCH under-covers consistently ($\approx$0.83–0.85) while being sharper than the calibrated methods, consistent with its conditional-Gaussian specification being slow to adapt to asymmetric tails during stress.
-Crisis-window behavior.
-Figure 1 (SP500, Feb–Apr 2020) shows how each method behaves during the shock.
-The red dots mark days where the realized return falls outside the reported 95% prediction interval (miscoverage).
-TCP/TCP–RM bands widen abruptly into the March spike and contract as conditions stabilize, whereas QR remains comparatively tight and under-covers around extremes.
-GARCH adapts smoothly via its volatility recursion, and Historical Simulation reacts with a lag due to its long lookback.
-Overall, all methods temporarily under-cover at the break, but TCP/TCP–RM inflate bands right when volatility jumps and recover calibration thereafter.
-Analogous crisis-window panels for BTC-USD and Gold are reported in Appendix B.
+TCP–RM’s coverage and width are essentially identical to TCP across assets (coverage deltas $\leq 0.0007$; width deltas $\leq 0.01$), indicating that the additional Robbins–Monro offset has minimal incremental effect at the chosen $(w,\gamma_{0})$ settings.
+GARCH under-covers consistently (about 82–85% coverage) while being sharper than the calibrated methods, consistent with its conditional-Gaussian specification being slow to adapt to asymmetric tails during stress.
+Crisis-window coverage.
+To stress-test the models, we also examine a 55-day “COVID crisis” window from February–April 2020 for each series.
+For this sub-period we compute the empirical coverage and average width of the 95% prediction intervals.
+Table 2 reports the results for the S&P 500; the corresponding tables for BTC-USD and Gold are given in Appendix B.
+For the S&P 500, TCP and TCP–RM both achieve 80% coverage with very similar average widths (11.94 and 11.95, respectively).
+GARCH attains the same coverage (80%) with a narrower interval (average width 9.66), while the rolling quantile regression and historical benchmark produce even sharper intervals (average widths 9.13 and 5.91) at the cost of substantially lower coverage (65.45% each).
+ACI is the most aggressive, with the narrowest intervals (average width 4.63) but only 56.36% coverage.
+Thus, during the sharpest part of the COVID shock, all methods under-cover, but TCP and TCP-RM remain competitive with GARCH in terms of the calibration-sharpness trade-off, while clearly dominating ACI and the nonparametric baselines in coverage.
+Table 2: Crisis-window empirical coverage and average width of 95% prediction intervals for the S&P 500 during the COVID shock (Feb–Apr 2020, $n=55$ days).
+Model Coverage Avg. width $n$
+TCP 0.8000 11.9371 55
+TCP-RM 0.8000 11.9511 55
+QR (rolling) 0.6545 9.1315 55
+GARCH 0.8000 9.6571 55
+Hist 0.6545 5.9058 55
+ACI 0.5636 4.6307 55
+TCP versus TCP–RM.
+Figure 2 overlays TCP and TCP–RM intervals with the realized returns during the COVID window for the S&P 500; analogous plots for BTC-USD and Gold are provided in Appendix C.
+The bands for TCP and TCP–RM are almost indistinguishable in all three series, and this is confirmed quantitatively.
+Over the COVID window, TCP and TCP–RM have identical empirical coverage for each asset (0.8000 for S&P 500, 0.8909 for BTC-USD, and 0.8727 for Gold), and their average interval widths differ only at the third decimal place: 11.94 vs 11.95 for S&P 500, 24.28 vs 24.27 for BTC-USD, and 6.81 vs 6.81 for Gold.
+Over the full sample, the two methods also remain nearly identical: for S&P 500, coverage is 0.9523 for TCP and 0.9530 for TCP–RM with average widths 5.2097 vs 5.2171 and interval scores 7.1168 vs 7.1157; for BTC-USD, both methods achieve coverage 0.9537 with widths 20.8879 vs 20.8859 and interval scores 27.9846 vs 27.9851; for Gold, both achieve coverage 0.9427 with widths 4.7368 vs 4.7438 and interval scores 6.3246 vs 6.3246.
+In other words, the Robbins–Monro adaptation layer leaves the marginal predictive performance of TCP essentially unchanged while preserving its good calibration properties.
+Figure 2: S&P 500: TCP vs TCP–RM 95% prediction intervals during the COVID crisis window (Feb–Apr 2020).
 Implementation note (comparability).
-All learning-based methods use the same feature set.
-TCP/TCP-RM are evaluated out-of-sample with a rolling window of 252 points split into 192 training and 60 calibration points.
-Rolling QR is also evaluated out-of-sample using a 192-point trailing training window.
-The slight differences in the “Predictions” counts (e.g., 1508 for QR vs.
-1448 for TCP/TCP-RM) reflect different warm-up requirements (QR does not reserve a calibration slice).
-Aligning the evaluation horizon to the longest warm-up yields qualitatively identical conclusions: QR is sharp but under-calibrated; TCP/TCP-RM are near-nominal and, in this evaluation, wider than Historical Simulation.
-6.1 Hyperparameter Sensitivity (SP500, TCP-RM)
-We stress–test the adaptive layer by varying the rolling window $w$ and the initial step size $\gamma_{0}$ for TCP-RM(Table 2).
-For comparability we fix the calibration slice at $|\mathcal{C}|=40$ and sweep $w\in\{100,252,500\}$ and $\gamma_{0}\in\{0.005,0.01,0.05\}$ on SP500.
-Results for BTC-USD and Gold exhibit the same pattern and are reported in Appendix D.
-Table 2: TCP-RM sensitivity on S&P 500 with $|\mathcal{C}|=40$ (target 95%).
-Hyperparameters Performance
-Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
-100 0.005 0.9575 5.3052
-100 0.010 0.9575 5.3066
-100 0.050 0.9581 5.3172
-252 0.005 0.9579 5.1168
-252 0.010 0.9579 5.1191
-252 0.050 0.9586 5.1314
-500 0.005 0.9525 5.4135
-500 0.010 0.9525 5.4319
-500 0.050 0.9567 5.5348
-What we learn.
-(i) Across these settings, w=252w{=}252 yields the narrowest intervals (width $\approx 5.12$), while w=100w{=}100 and w=500w{=}500 are wider ($\approx 5.31$ and 5.415.415.535.53, respectively), with coverage remaining near or just above the 95% target for $w\in\{100,252\}$ and slightly lower for some w=500w{=}500 configurations.
-(ii) At a fixed $w$, increasing $\gamma_{0}$ produces only minor changes: coverage nudges up by at most a few tenths of a percentage point with a negligible width increase (e.g., at w=252w{=}252, $\gamma_{0}{=}0.050$ lifts coverage from $0.9579$ to $0.9586$ with a $\sim 0.012$ increase in width; at w=500w{=}500, $\gamma_{0}{=}0.050$ raises coverage from $0.9525$ to $0.9567$ while widening intervals modestly).
-Takeaway.
-TCP–RM is robust over $(w,\gamma_{0})$ in this range.
-For the main experiments, w=252w{=}252 with $\gamma_{0}{=}0.01$ offers a good balance (coverage $\approx 0.9579$, width $\approx 5.1191$).
-If one prefers higher coverage at w=500w{=}500, a slightly larger $\gamma_{0}$ (e.g., $0.05$) helps, at the cost of modestly wider bands.
-We also perform sensitivity analysis on TCP. The results are given in Appendix E.
-6.2 Visualization for TCP vs. TCP-RM on S&P 500 (95% intervals)
-In this section, we present the TCP vs TCP-RM trace plot for S&P 500.
-Figure 2: TCP vs.
-TCP-RM on S&P 500 (95% intervals).
-Top: daily returns with two‐sided prediction bands for TCP (blue) and TCP-RM (purple).
-Middle: 30-day rolling empirical coverage; the dashed line marks the 95% target.
-Bottom: evolution of conformal thresholds: split‐conformal level $C_{\text{split}}$, online Robbins–Monro offset $C_{\text{RM}}$, and the effective total $C_{\text{eff}}=C_{\text{split}}+C_{\text{RM}}$.
-During the COVID-19 crash the bands widen sharply, the rolling coverage briefly departs from the 95% target and then re-enters the band, and $C_{\text{eff}}$ spikes before decaying as conditions stabilize.
-In Fig.
-2, TCP and TCP-RM behave similarly in tranquil periods, but during stress the effective threshold $C_{\text{eff}}$ (bottom panel) rises quickly, producing wider bands (top) and steering the 30-day rolling coverage back toward the 95% target (middle).
-The Robbins–Monro offset $C_{\text{RM}}$ remains near zero except around regime shifts, where it provides the extra adjustment that closes the gap to target coverage.
-This visual evidence complements Table 1: on S&P 500, TCP-RM attains $\approx$95% coverage with only a marginal width increase relative to TCP.
+All learning-based methods (TCP, TCP-RM, ACI, QR, QR-Linear) use the same feature set.
+TCP/TCP-RM and ACI are evaluated out-of-sample with a rolling window of 252 points split into 192 training and 60 calibration points.
+Rolling QR and QR-Linear are also evaluated out-of-sample using a 192-point trailing training window without a separate calibration slice.
+The differences in the “Predictions” counts (e.g., 1508 for QR/QR-Linear vs.
+1448 for TCP/TCP-RM/ACI, 1468 for Hist, 1670 for GARCH) reflect these distinct warm-up requirements.
+Aligning the evaluation horizon to the longest warm-up yields qualitatively identical conclusions: tree-based QR remains sharp but under-calibrated; QR-Linear and ACI partially trade width for better coverage yet stay sub-nominal; and TCP/TCP-RM are near-nominal and, in this evaluation, wider than Historical Simulation.
+6.1 Backtests of tail coverage
+To complement marginal coverage and interval width, we perform standard VaR-style backtests on the 5% lower tail using the Kupiec (Kupiec1995) unconditional coverage test (LR_uc), the Christoffersen (Christoffersen1998) independence test (LR_ind), and their joint conditional coverage test (LR_cc).
+Table 3 reports the likelihood-ratio statistics, $p$-values, the total number of exceedances (days where the realized return falls below the nominal 5% lower bound), and the number of out-of-sample predictions $n$ for the S&P 500.
+Table 3: VaR-style backtests on the 5% lower tail for S&P 500.
+Model LR_uc p_uc LR_ind p_ind LR_cc p_cc Exceed. $n$
+TCP 0.170625 0.679557 2.022260 0.155008 2.192885 0.334057 69 1448
+TCP-RM 0.287050 0.592117 2.203339 0.137712 2.490389 0.287885 68 1448
+ACI 127.325197 0 7.171877 0.007406 134.497073 0 183 1448
+QR 280.342540 0 1.404329 0.236000 281.746869 0 253 1508
+QR-Linear 100.761827 0 10.714099 0.001063 111.475925 0 174 1508
+GARCH 334.481845 0 0.022234 0.881466 334.504079 0 289 1670
+Hist 9.827570 0.001719 22.904448 0.000002 32.732018 0 101 1468
+On S&P 500, TCP and TCP-RM are the only methods that pass all three backtests at conventional significance levels: their LR_uc, LR_ind, and LR_cc $p$-values stay well above 0.05, and the number of exceedances (68–69) is close to the 5% target ($\approx 0.05\times 1448$).
+By contrast, ACI, QR, QR-Linear, GARCH, and Historical Simulation are all strongly rejected by at least one of the tests: their LR_uc and LR_cc statistics are very large with $p$-values numerically equal to zero (e.g., QR: 253 exceedances; GARCH: 289; Hist: 101).
+This confirms that, on equities, only TCP/TCP-RM deliver statistically acceptable 95% tail coverage.
+The corresponding backtests for BTC-USD and Gold show the same qualitative pattern and are reported in Appendix D.
+We also perform sensitivity analyses for TCP-RM and TCP, with results reported in Appendix F and Appendix G. TCP vs. TCP-RM trace plots are presented in Appendix H.
 7.  CONCLUSION AND FUTURE WORK
-This paper introduced Temporal Conformal Prediction (TCP), a simple and effective framework for calibrated prediction intervals in nonstationary time series.
-TCP combines a rolling, exchangeability‐aware split–conformal layer with a modern quantile forecaster, and the TCP–RM variant adds a single Robbins–Monro offset that adjusts the conformal threshold online using observed coverage errors.
-Across three asset classes, TCP/TCP–RM delivers a strong calibration–sharpness balance: coverage is close to the 95% target; in our evaluation, the intervals are wider than Historical Simulation, reflecting stricter calibration during volatile periods.
-In contrast, the rolling QR baseline attains the narrowest bands but undershoots the target coverage, especially in turbulent regimes.
-Crisis‐window visualizations around the COVID‐19 crash illustrate TCP’s intended adaptiveness: intervals widen rapidly as volatility spikes and contract as conditions stabilize, with red dots indicating miscoverage events.
-A targeted sensitivity analysis shows robustness to reasonable choices of rolling window and Robbins–Monro step size, and suggests default settings that are stable in practice.
-Limitations.
-Like any sequential method, TCP can temporarily under‐cover during abrupt regime shifts before its calibration catches up.
-Performance also inherits the quality of the base quantile forecaster and the choice of windowing; extremely heavy tails or structural breaks can stress any fixed feature set.
-Finally, while TCP–RM improves calibration at negligible width cost on average, very aggressive step sizes can overcorrect.
-Practical implications.
-By providing distribution‐free intervals that adapt in time, TCP aligns naturally with backtesting and governance needs in risk management (coverage targets, exception tracking, and scenario robustness).
-The framework is lightweight (single threshold) and model‐agnostic (plugs into any quantile forecaster), making deployment straightforward in production pipelines.
-7.1 Future Work
-Several extensions follow naturally:
-  1. 1.
-    Multivariate and portfolio TCP. Generalize to joint/conditional quantiles for baskets of assets; couple TCP with copula or factor models to control portfolio‐level coverage.
-  2. 2.
-    Conditional and localized coverage. Target coverage conditional on covariates (e.g., volatility state) and explore adaptive nonconformity scores that emphasize tail risk or signed errors.
-  3. 3.
-    Richer forecasters and features.
-Swap the base learner for neural quantile models or boosted variants with realized‐volatility inputs (intraday RV), order‐flow, or macro features; assess benefits for sharpness at fixed coverage.
-  4. 4.
-    Change–point aware adaptation. Trigger temporary step–size boosts or threshold projections when a regime break is detected to shorten the transient under/over‐shoot in coverage.
-  5. 5.
-    Heavy‐tail robustness. Investigate EVT‐inspired nonconformity scores or asymmetric penalties to better stabilize performance under extreme shocks.
-  6. 6.
-    Decision integration. Map TCP intervals to VaR/ES, margin, and capital rules; evaluate utility/capital efficiency vs. conservative baselines in realistic backtests.
-  7. 7.
-    Beyond finance. TCP’s adaptiveness is broadly applicable wherever nonstationarity and calibrated uncertainty matter:
-    • Biostatistics/medicine: ICU vitals monitoring and glucose‐range forecasting with patient–specific adaptation; adaptive dosing intervals under drift.
-    • Epidemiology: Calibrated intervals for nowcasts/forecasts of incidence and $R_{t}$ that widen under reporting delays or testing-policy shifts, plus adaptive risk bands for hospitalization/ICU load as transmission regimes change.
-    • Genetics/Genomics: Uncertainty bands for polygenic risk scores and eQTL effect sizes that remain calibrated under population/covariate shift; interval forecasts for gene expression (bulk or single-cell) with batch effects and sequencing-depth drift; calibrated pathogenicity scores for rare variants.
-    • Sports analytics: Game-state–aware intervals for win probability, expected goals/points, and player performance that adapt to tempo and lineup changes; injury-risk ranges for load management with rapid recalibration after shocks (e.g., in-game injuries or schedule congestion).
-    • Renewable energy: Wind/solar generation intervals that expand/contract with weather regimes to aid grid balancing.
-    • Operations/marketing: Demand and lead‐time intervals for inventory control; calibrated lift ranges in marketing‐mix models under campaign drift.
-    • Transportation and climate: Travel‐time intervals under traffic regime shifts; local hazard/risk bands for extreme weather.
-In sum, TCP offers a compact, distribution‐free, and adaptive route to calibrated uncertainty quantification in the presence of temporal drift.
-Its empirical performance, visual adaptiveness in crisis windows, and robustness to sensible hyperparameters suggest it is a practical default for risk forecasting and a promising building block for broader sequential decision systems.
+This paper introduced Temporal Conformal Prediction (TCP), a simple framework for calibrated prediction intervals in nonstationary time series.
+TCP combines a rolling, exchangeability-aware split-conformal layer with a modern quantile forecaster, while the TCP–RM variant adds a single Robbins–Monro offset that adjusts the conformal threshold online using observed coverage errors.
+Across three asset classes, TCP and TCP–RM achieve a strong calibration–sharpness balance: coverage stays close to the 95% target and the intervals are wider than Historical Simulation, reflecting stricter calibration in volatile periods.
+The rolling tree-based QR model and the classical linear baseline (QR-Linear) produce the narrowest bands but undershoot the target coverage, especially in turbulent regimes, and the adaptive conformal baseline (ACI) only partially corrects this undercoverage.
+In our study, TCP and TCP-RM are the only methods that simultaneously passes standard VaR backtests and maintains competitive sharpness on S&P 500, BTC-USD, and Gold.
+Like any sequential method, TCP can temporarily under-cover during abrupt regime shifts before calibration catches up.
+Performance depends on the quality of the base quantile forecaster and the windowing scheme, and very heavy tails or structural breaks can stress any fixed feature set.
+In our experiments, TCP–RM leaves average calibration and sharpness essentially unchanged at the chosen hyperparameters, but its Robbins–Monro layer provides a principled online mechanism that can correct coverage drift when step sizes are tuned more aggressively; overly aggressive step sizes can, however, overcorrect.
+On the positive side, TCP provides distribution-free, time-adaptive intervals that align naturally with backtesting and governance needs in risk management, and its single-threshold, model-agnostic design makes deployment straightforward in production pipelines.
+Several extensions follow naturally.
+On the modeling side, it is natural to develop multivariate and portfolio versions of TCP, target conditional or localized coverage (for example by volatility state), explore richer quantile forecasters and feature sets, design change-point-aware adaptations that shorten transients after regime breaks, and study heavy-tail robustness through EVT-inspired or asymmetric nonconformity scores.
+On the application side, TCP intervals can be mapped directly to VaR/ES, margin, and capital rules and evaluated in realistic backtests for utility and capital efficiency.
+Beyond finance, the same adaptive, distribution-free mechanism could support calibrated uncertainty in domains such as biostatistics and medicine (for example ICU vitals or glucose-range forecasting), epidemiology, genetics and genomics, sports analytics, renewable energy, and operations or transportation planning wherever nonstationarity and reliable uncertainty quantification are central.
+In sum, TCP offers a compact, distribution-free, and adaptive route to calibrated uncertainty quantification under temporal drift.
+Its empirical performance, visual adaptiveness in crisis windows, and robustness to sensible hyperparameters suggest that it is a practical default for risk forecasting and a promising building block for broader sequential decision systems.
+Code Availability
 Appendix A MATHEMATICAL PROOFS
@@
 Equivalently, $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{r_{t}\in[\ell_{t},u_{t}]\}\to 1-\alpha$ a.s., which completes the proof.
-Appendix B SUPPLEMENTARY VISUALIZATIONS
+Appendix B ADDITIONAL CRISIS-WINDOW RESULTS
+Table 4 and Table 5 report the COVID crisis-window coverage and widths for BTC-USD and Gold.
+Table 4: Crisis-window empirical coverage and average width of 95% prediction intervals for BTC-USD during COVID (Feb–Apr 2020, $n=55$ days).
+Model Coverage Avg. width $n$
+TCP 0.8909 24.2814 55
+TCP-RM 0.8909 24.2729 55
+QR (rolling) 0.8909 17.4407 55
+GARCH 0.8545 17.9290 55
+Hist 0.9091 19.8499 55
+ACI 0.8727 20.2050 55
+Table 5: Crisis-window empirical coverage and average width of 95% prediction intervals for Gold during COVID (Feb–Apr 2020, $n=55$ days).
+Model Coverage Avg. width $n$
+TCP 0.8727 6.8087 55
+TCP-RM 0.8727 6.8119 55
+QR (rolling) 0.7455 4.0744 55
+GARCH 0.8000 4.9112 55
+Hist 0.6909 3.9957 55
+ACI 0.5818 2.6643 55
+Appendix C ADDITIONAL TCP VS TCP-RM CRISIS-WINDOW PLOTS
+Figure 3 and Figure 4 show the TCP and TCP-RM 95% prediction intervals alongside returns for BTC-USD and Gold during the COVID crisis window (Feb–Apr 2020).
+Figure 3: BTC-USD: TCP vs TCP-RM 95% prediction intervals during the COVID crisis window (Feb–Apr 2020).
+Figure 4: Gold: TCP vs TCP-RM 95% prediction intervals during the COVID crisis window (Feb–Apr 2020).
+Appendix D ADDITIONAL BACKTESTS
+Table 6 reports the same Kupiec and Christoffersen 5% lower-tail backtests for BTC-USD and Gold.
+The qualitative pattern matches the S&P 500 results in the main text: TCP and TCP-RM are the only models that are not rejected by any of the three tests on either asset, while ACI, QR, QR-Linear, GARCH, and Historical Simulation accumulate far too many exceptions and fail at least one of the diagnostics.
+Table 6: VaR-style backtests on the 5% lower tail for BTC-USD and Gold.
+Asset Model LR_uc p_uc LR_ind p_ind LR_cc p_cc Exceed. $n$
+BTC-USD TCP 0.434337 0.509868 1.094155 0.295552 1.528492 0.465685 67 1448
+TCP-RM 0.434337 0.509868 1.094155 0.295552 1.528492 0.465685 67 1448
+ACI 89.465919 0 11.093370 0.000866 100.559289 0 163 1448
+QR 192.386083 0 0.868073 0.351489 193.254155 0 218 1508
+QR-Linear 73.429698 0 5.932611 0.014863 79.362309 0 158 1508
+GARCH 221.469896 0 1.342326 0.246624 222.812222 0 245 1670
+Hist 1.023574 0.311673 9.776309 0.001768 10.799882 0.004517 82 1468
+Gold TCP 1.563187 0.211199 0.338333 0.560793 1.901519 0.386447 83 1448
+TCP-RM 1.563187 0.211199 0.338333 0.560793 1.901519 0.386447 83 1448
+ACI 280.191360 0 2.694226 0.100712 282.885586 0 247 1448
+QR 288.442776 0 1.452719 0.228092 289.895494 0 256 1508
+QR-Linear 78.288019 0 13.470197 0.000242 91.758217 0 161 1508
+GARCH 291.399454 0 1.952267 0.162343 293.351721 0 273 1670
+Hist 8.513620 0.003525 2.757010 0.096829 11.270630 0.003570 99 1468
+Appendix E ADDITIONAL VISUALIZATIONS OF THE PREDICTION INTERVALS
 This section provides the prediction interval visualizations for BTC-USD (5) and Gold (6) during the COVID-19 market crash (February-April 2020).
-Figure 3: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for BTC-USD daily returns during the COVID-19 market crash (February-April 2020).
-Figure 4: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for Gold daily returns during the COVID-19 market crash (February-April 2020).
-Appendix C VISUALIZATIONS OF THE TCP VS TCP-RM TRACE PLOTS
-This section provides the visualizations of the TCP vs TCP-RM trace plots for BTC-USD (Figure 5) and Gold (Figure 6), complementing the main visualization for the S&P 500 presented in the results section.
-Figure 5: TCP vs.
-TCP-RM on BTC-USD (95% intervals).
-Same layout as Fig.
-2.
-Large, frequent shocks lead to sustained but controlled elevations in $C_{\text{eff}}$, keeping rolling coverage close to 95% while allowing intervals to shrink as volatility recedes.
-Figure 6: TCP vs.
-TCP-RM on Gold (95% intervals).
-Same layout as Fig.
-2.
-Spikes in $C_{\text{eff}}$ are concentrated around crisis windows; elsewhere, $C_{\text{RM}}$ stays near zero and the method behaves like standard split-conformal TCP.
-Appendix D SUPPLEMENTARY SENSITIVITY ANALYSIS
-This section provides the full sensitivity analysis results for BTC-USD (Table 3) and Gold (Table 4), complementing the main analysis on the S&P 500 presented in the results section.
-The findings confirm that the TCP-RM framework is robust and performs consistently across different asset classes and hyperparameter settings.
-For BTC-USD, widths decrease as $w$ grows (tightest at w=500w{=}500) while coverage remains near 95%–96%; $\gamma_{0}$ has negligible impact within a fixed $w$.
-For Gold, widths are smallest at w=100w{=}100 and increase with $w$; coverage hovers around the nominal level; $\gamma_{0}$ effects are minor at fixed $w$.
+Figure 5: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist, ACI) for BTC-USD daily returns during the COVID-19 market crash (February-April 2020).
+Figure 6: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist, ACI) for Gold daily returns during the COVID-19 market crash (February-April 2020).
+Appendix F SENSITIVITY ANALYSIS FOR TCP-RM
+This section reports the full hyperparameter sensitivity analysis for TCP-RM across the S&P 500, BTC-USD, and Gold.
+For each asset we vary the rolling window size $w\in\{100,252,500\}$ and the initial Robbins-Monro step size $\gamma_{0}\in\{0.005,0.01,0.05\}$ while keeping the calibration slice fixed at $|\mathcal{C}|=40$.
+For every configuration, we report empirical coverage of the nominal 95% prediction intervals and average interval width (Tables 7,8, and 9).
+Table 7: TCP-RM sensitivity on S&P 500 with $|\mathcal{C}|=40$ (target 95%).
+Hyperparameters Performance
+Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
+100 0.005 0.9575 5.3052
+100 0.010 0.9575 5.3066
+100 0.050 0.9581 5.3172
+252 0.005 0.9579 5.1168
+252 0.010 0.9579 5.1191
+252 0.050 0.9586 5.1314
+500 0.005 0.9525 5.4135
+500 0.010 0.9525 5.4319
+500 0.050 0.9567 5.5348
 Table 8: TCP-RM sensitivity on BTC-USD with $|\mathcal{C}|=40$.
@@
 500 0.050 0.9542 5.2031
-Appendix E SUPPLEMENTARY SENSITIVITY ANALYSIS ON TCP (NO RM)
-For completeness, in Tables 5, 6 and 7, we also report sensitivity for TCP (split-conformal only, no Robbins–Monro offset). We vary $w\in\{100,252,500\}$ with $|\mathcal{C}|=40$ for each asset.
+Summary.
+On S&P 500, w=252w{=}252 yields the narrowest intervals (width $\approx 5.12$), with w=100w{=}100 and w=500w{=}500 producing slightly wider bands and coverage close to or just above the 95% target, except for mild undercoverage in some w=500w{=}500 settings; this motivates our choice w=252w{=}252, $\gamma_{0}{=}0.01$ in the main experiments.
+For BTC-USD, interval widths shrink as $w$ increases, with the tightest bands at w=500w{=}500, while coverage remains stably around 95–96% across all configurations.
+For Gold, widths are smallest at w=100w{=}100 and grow with $w$, and coverage stays near the nominal level throughout.
+Across all three assets, TCP-RM is therefore stable over the $(w,\gamma_{0})$ grid considered here: the window length $w$ controls the sharpness–stability tradeoff, whereas the specific choice of $\gamma_{0}$ within this range has only a minor effect on coverage and width.
+Appendix G SENSITIVITY ANALYSIS ON TCP (NO RM)
+For completeness, in Tables 10, 11 and 12, we also report sensitivity for TCP (split-conformal only, no Robbins–Monro offset). We vary $w\in\{100,252,500\}$ with $|\mathcal{C}|=40$ for each asset.
 Table 10: TCP sensitivity on S&P 500 with $|\mathcal{C}|=40$.
@@
 Without the adaptive offset, coverage can dip slightly below 95% in some settings, underscoring the benefit of the online calibration layer.
+Appendix H TCP VS. TCP-RM TRACE PLOTS
+This section provides TCP and TCP-RM trace plots for all three assets (Figures 7, 8 and 9.
+Each panel shows daily returns, the corresponding 95% prediction intervals, a rolling coverage diagnostic, and the evolution of the conformal thresholds.
+Figure 7: TCP vs.
+TCP-RM on S&P 500 (95% intervals).
+Top: daily returns with two-sided prediction bands for TCP (blue) and TCP-RM (purple).
+Middle: 30-day rolling empirical coverage; the dashed line marks the 95% target.
+Bottom: evolution of conformal thresholds: split-conformal level $C_{\text{split}}$, Robbins–Monro offset $C_{\text{RM}}$, and effective total $C_{\text{eff}}=C_{\text{split}}+C_{\text{RM}}$.
+During the COVID-19 crash the bands widen sharply, the rolling coverage briefly moves away from the 95% target and then returns, and $C_{\text{eff}}$ spikes before decaying as conditions stabilize.
+Figure 8: TCP vs.
+TCP-RM on BTC-USD (95% intervals).
+Same layout as Figure 7.
+Large and frequent shocks lead to sustained but controlled elevations in $C_{\text{eff}}$, which keeps rolling coverage close to 95% while allowing intervals to shrink as volatility recedes.
+Figure 9: TCP vs.
+TCP-RM on Gold (95% intervals).
+Same layout as Figure 7.
+Spikes in $C_{\text{eff}}$ are concentrated around crisis windows; away from those periods $C_{\text{RM}}$ stays close to zero and the method behaves like standard split-conformal TCP.

### v6

@@
 We propose Temporal Conformal Prediction (TCP), a distribution-free framework for constructing well-calibrated prediction intervals in nonstationary time series.
-TCP couples a modern quantile forecaster with a split-conformal calibration layer on a rolling window and, in its TCP-RM variant, augments the conformal threshold with a single online Robbins-Monro (RM) offset to steer coverage toward a target level in real time.
-We benchmark TCP against GARCH, Historical Simulation, a rolling tree-based Quantile Regression (QR) model, a classical linear quantile regression baseline (QR-Linear), and an adaptive conformal method (ACI) across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold).
-Three results are consistent across assets.
-First, both QR and QR-Linear yield the sharpest intervals but are materially under-calibrated, and even ACI remains below the nominal 95% target in our full-sample backtests.
-Second, TCP (and TCP-RM) achieves near-nominal coverage across assets, with intervals that are wider than Historical Simulation in this evaluation (e.g., S&P 500: 5.21 vs.
+TCP couples a modern quantile forecaster with a rolling split-conformal calibration layer; its TCP-RM variant adds an online Robbins-Monro offset to steer coverage in real time.
+We benchmark TCP against GARCH, Historical Simulation, Quantile Regression (QR), linear QR, and Adaptive Conformal Inference (ACI) across S&P 500, Bitcoin, and Gold.
+Three results are consistent.
+First, QR baselines yield the sharpest intervals but are materially under-calibrated; even ACI remains below the 95% target.
+Second, TCP achieves near-nominal coverage, yielding intervals slightly wider than Historical Simulation (e.g., S&P 500: 5.21 vs.
 5.06).
-Third, the RM update changes calibration and width only marginally at our default hyperparameters.
-Crisis-window visualizations around March 2020 show TCP/TCP-RM expanding and then contracting their interval bands promptly as volatility spikes and recedes, with red dots marking days where realized returns fall outside the reported 95% interval (miscoverage).
-A sensitivity study confirms robustness to window size and step-size choices.
-Overall, TCP provides a practical, theoretically grounded solution to calibrated uncertainty quantification under distribution shift, bridging statistical inference and machine learning for risk forecasting.
+Third, the RM update changes calibration only marginally at default hyperparameters.
+Crisis-window visualizations (March 2020) show TCP promptly expanding and contracting intervals as volatility spikes.
+A sensitivity study confirms robustness to hyperparameters.
+Overall, TCP bridges statistical inference and machine learning, providing a practical solution for calibrated risk forecasting under distribution shift.
 Keywords: Temporal Conformal Prediction, Quantile Regression, Machine Learning, Statistical Learning, Value-at-Risk, Financial Time Series, Risk Management.
-MSC 2020 Subject Classification: 62G08, 62M10, 62P05, 91G70, 68T05
 1.  Introduction
-Financial risk estimation is more than a regulatory checkbox; it is foundational for market stability and investor confidence (Markowitz1952; Morgan1996).
+Financial risk estimation is more than a regulatory checkbox; it is foundational for market stability and investor confidence (Markowitz, 1952; J.P.
+Morgan/Reuters, 1996).
 Yet, when markets enter turbulent regimes, traditional risk models often fall short.
-Early implementations of Value-at-Risk (VaR) were predominantly parametric, often assuming (conditional) normality (Jorion2007; Dowd1998), assumptions that can systematically understate tail risk.
-Coherent alternatives such as Conditional Value-at-Risk (CVaR) were introduced to address some of these limitations (Rockafellar2000).
-Events like the 2008 global financial crisis (Brunnermeier2009) and the March 2020 stock-market crash (Mazur2021) exposed the fragility of parametric assumptions, especially in the tails.
+Early implementations of Value-at-Risk (VaR) were predominantly parametric, often assuming (conditional) normality (Jorion, 2007; Dowd, 1998), assumptions that can systematically understate tail risk.
+Coherent alternatives such as Conditional Value-at-Risk (CVaR) were introduced to address some of these limitations (Rockafellar and Uryasev, 2000).
+Events like the 2008 global financial crisis (Brunnermeier, 2009) and the March 2020 stock-market crash (Mazur et al., 2021) exposed the fragility of parametric assumptions, especially in the tails.
 The core challenge is that financial returns violate the independent and identically distributed (i.i.d.) assumption that underpins many statistical learning techniques.
-Real-world returns are heteroskedastic, exhibit regime shifts, and often exhibit heavy tails (Cont2001).
-GARCH-type models (Engle1982; Bollerslev1986) can capture some volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
+Real-world returns are heteroskedastic, exhibit regime shifts, and often exhibit heavy tails (Cont, 2001).
+GARCH-type models (Engle, 1982; Bollerslev, 1986) can capture some volatility clustering, but they still depend on pre-specified distributions and lack reliable coverage guarantees for interval forecasts.
 In recent years, conformal prediction has gained traction as a powerful distribution-free framework for quantifying uncertainty.
-Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk2005).
+Classical conformal methods provide finite-sample coverage guarantees without assuming a particular data distribution (Vovk et al., 2005).
 However, these guarantees rest on exchangeability, a condition often violated by time series with temporal dependencies.
-Several works extend conformal methods to sequential or temporally dependent settings (Xu2021; Stankeviciute2021; Gibbs2021), but a robust and practical solution for financial markets remains an open challenge.
+Several works extend conformal methods to sequential or temporally dependent settings (Xu and Xie, 2021; Stankeviciute et al., 2021; Gibbs and Candès, 2021), but a robust and practical solution for financial markets remains an open challenge.
 This paper closes that gap.
@@
 Traditional Financial Risk Models.
-Quantitative risk management traces its roots to portfolio theory (Markowitz1952) and was formalized around the VaR concept with J.P.
-Morgan’s RiskMetrics (Morgan1996).
-In practice, three VaR paradigms have dominated: parametric approaches (often assuming conditional normality), non-parametric Historical Simulation, and Monte Carlo methods (Jorion2007; Dowd1998; Glasserman2004).
-Beyond VaR, CVaR provided a coherent alternative for tail-risk control (Rockafellar2000).
-However, many traditional approaches underperform in crisis backtests (Kupiec1995; Christoffersen1998), highlighting a blind spot when tail events dominate.
+Quantitative risk management traces its roots to portfolio theory (Markowitz, 1952) and was formalized around the VaR concept with J.P.
+Morgan’s RiskMetrics (J.P.
+Morgan/Reuters, 1996).
+In practice, three VaR paradigms have dominated: parametric approaches (often assuming conditional normality), non-parametric Historical Simulation, and Monte Carlo methods (Jorion, 2007; Dowd, 1998; Glasserman, 2004).
+Beyond VaR, CVaR provided a coherent alternative for tail-risk control (Rockafellar and Uryasev, 2000).
+However, many traditional approaches underperform in crisis backtests (Kupiec, 1995; Christoffersen, 1998), highlighting a blind spot when tail events dominate.
 Advanced Time-Series and Machine-Learning Models.
-To better capture time-varying variance and volatility clustering, the GARCH family became foundational (Engle1982; Bollerslev1986); asymmetric variants such as EGARCH and GJR-GARCH address leverage and sign effects (Nelson1991; Glosten1993).
-For heavy-tailed markets, extreme value theory (EVT) offers principled tail modeling (Pickands1975; McNeil2000), and VaR under heavy tails raises specific concerns (Danielsson2000).
-On the machine-learning side, quantile regression directly targets conditional quantiles (Koenker1978), with ensemble and neural variants such as quantile regression forests (Meinshausen2006) and quantile networks (Taylor2000); deep LSTMs capture complex temporal dependencies (Fischer2018).
-We also emphasize the computational efficiency of gradient-boosted trees via LightGBM for quantile forecasting (Ke2017).
+To better capture time-varying variance and volatility clustering, the GARCH family became foundational (Engle, 1982; Bollerslev, 1986); asymmetric variants such as EGARCH and GJR-GARCH address leverage and sign effects (Nelson, 1991; Glosten et al., 1993).
+For heavy-tailed markets, extreme value theory (EVT) offers principled tail modeling (Pickands, 1975; McNeil and Frey, 2000), and VaR under heavy tails raises specific concerns (Danielsson and de Vries, 2000).
+On the machine-learning side, quantile regression directly targets conditional quantiles (Koenker and Bassett, 1978), with ensemble and neural variants such as quantile regression forests (Meinshausen, 2006) and quantile networks (Taylor, 2000); deep LSTMs capture complex temporal dependencies (Fischer and Krauss, 2018).
+We also emphasize the computational efficiency of gradient-boosted trees via LightGBM for quantile forecasting (Ke et al., 2017).
 Conformal Prediction Theory.
-Conformal prediction provides distribution-free, finite-sample prediction sets under exchangeability (Vovk2005).
-Recent advances relax these assumptions for sequential or non-stationary settings through adaptive calibration and time-series‐aware procedures (Gibbs2021; Xu2021; Stankeviciute2021).
+Conformal prediction provides distribution-free, finite-sample prediction sets under exchangeability (Vovk et al., 2005).
+Recent advances relax these assumptions for sequential or non-stationary settings through adaptive calibration and time-series‐aware procedures (Gibbs and Candès, 2021; Xu and Xie, 2021; Stankeviciute et al., 2021).
 Our work builds on these foundations with a practical adaptive conformal framework tailored to financial risk forecasting, combining local split-conformal calibration with an online Robbins–Monro offset for temporal adaptation.
@@
 Recent research has addressed non-stationary environments by adapting the conformal guarantee.
-Gibbs2021 and zaffran2022adaptive introduce Adaptive Conformal Inference (ACI), which updates the target miscoverage rate ($\alpha_{t}$) online based on past coverage errors.
+Gibbs and Candès (2021) and Zaffran et al.
+(2022) introduce Adaptive Conformal Inference (ACI), which updates the target miscoverage rate ($\alpha_{t}$) online based on past coverage errors.
 While effective, we argue that in heavy-tailed financial settings, oscillating the target probability can lead to unstable interval widths; TCP-RM instead adapts the conformal threshold ($C_{t}$) directly via a Robbins-Monro update, offering a stable, scalar lever for width control.
-Other approaches, such as CPTD (lin2022temporal), leverage cross-sectional exchangeability across a panel of time series to restore validity.
+Other approaches, such as CPTD (Lin et al., 2022), leverage cross-sectional exchangeability across a panel of time series to restore validity.
 In contrast, TCP is designed for the univariate setting (single asset) where cross-sectional pooling (e.g., mixing Bitcoin and Gold) is inappropriate due to distinct, idiosyncratic asset dynamics.
@@
 This section lays the theoretical groundwork for our proposed method.
-Our calibration mechanism is related in spirit to adaptive conformal procedures for nonstationary settings (zaffran2022adaptive; lin2022temporal) but differs in using a Robbins–Monro update on the conformal threshold while holding the nominal level fixed.
+Our calibration mechanism is related in spirit to adaptive conformal procedures for nonstationary settings (Zaffran et al., 2022; Lin et al., 2022) but differs in using a Robbins–Monro update on the conformal threshold while holding the nominal level fixed.
 We begin by formally defining the problem of prediction interval forecasting, then review the principles of classical conformal prediction and its limitations, and finally develop the core mathematical components of our Temporal Conformal Prediction (TCP) framework and its adaptive calibration mechanism.
@@
     Stochastic approximation framework.
-The Robbins-Monro update for the offset is analyzed using classical stochastic-approximation arguments under mixing and bounded-variance noise; our assumptions instantiate those conditions in this setting ((Kushner2003)).
+The Robbins-Monro update for the offset is analyzed using classical stochastic-approximation arguments under mixing and bounded-variance noise; our assumptions instantiate those conditions in this setting ((Kushner and Yin, 2003)).
   2. 2.
@@
     Assumption verification in practice.
-We empirically probe the mixing/independence assumptions using standard risk diagnostics (Kupiec1995 and Christoffersen1998 backtests); the lack of serial dependence in exceedances supports the validity of these conditions on our financial datasets.
+We empirically probe the mixing/independence assumptions using standard risk diagnostics (Kupiec (1995) and Christoffersen (1998) backtests); the lack of serial dependence in exceedances supports the validity of these conditions on our financial datasets.
 4.5 Proof of Asymptotic Calibration
@@
   3. 3.
-    Adaptive Conformal Inference (ACI): A rolling implementation of Gibbs2021 that uses the same window split as TCP ($w_{\mathrm{tr}}=192$ training, $w_{\mathrm{cal}}=60$ calibration).
+    Adaptive Conformal Inference (ACI): A rolling implementation of Gibbs and Candès (2021) that uses the same window split as TCP ($w_{\mathrm{tr}}=192$ training, $w_{\mathrm{cal}}=60$ calibration).
 However, instead of updating the conformal threshold $C_{t}$ (like TCP-RM), ACI updates the target miscoverage rate $\alpha_{t}$ online based on the most recent coverage error (step size $\eta=0.05$).
@@
 For each model and asset, we also perform standard 1-day-ahead VaR backtests on the 5% lower tail.
-Specifically, we report: (i) the unconditional coverage test of Kupiec1995 (LRuc), (ii) the independence test of Christoffersen1998 (LRind), and (iii) their sum $_{\mathrm{cc}}{=}LR$$_{\mathrm{uc}}{+}$LRind for conditional coverage.
+Specifically, we report: (i) the unconditional coverage test of Kupiec (1995) (LRuc), (ii) the independence test of Christoffersen (1998) (LRind), and (iii) their sum $_{\mathrm{cc}}{=}LR$$_{\mathrm{uc}}{+}$LRind for conditional coverage.
 Each test yields a likelihood-ratio statistic and a $p$-value, together with the number of exceedances $x$ (days where the realized return falls below the nominal 5% VaR) out of $n$ out-of-sample predictions.
@@
 6.1 Backtests of tail coverage
-To complement marginal coverage and interval width, we perform standard VaR-style backtests on the 5% lower tail using the Kupiec (Kupiec1995) unconditional coverage test (LR_uc), the Christoffersen (Christoffersen1998) independence test (LR_ind), and their joint conditional coverage test (LR_cc).
+To complement marginal coverage and interval width, we perform standard VaR-style backtests on the 5% lower tail using the Kupiec (Kupiec, 1995) unconditional coverage test (LR_uc), the Christoffersen (Christoffersen, 1998) independence test (LR_ind), and their joint conditional coverage test (LR_cc).
 Table 3 reports the likelihood-ratio statistics, $p$-values, the total number of exceedances (days where the realized return falls below the nominal 5% lower bound), and the number of out-of-sample predictions $n$ for the S&P 500.
@@
 The corresponding backtests for BTC-USD and Gold show the same qualitative pattern and are reported in Appendix D.
-We also perform sensitivity analyses for TCP-RM and TCP, with results reported in Appendix F and Appendix G. TCP vs. TCP-RM trace plots are presented in Appendix H.
+6.2 Sensitivity Analysis
+This section reports the full hyperparameter sensitivity analysis for TCP-RM across the S&P.
+We vary the rolling window size $w\in\{100,252,500\}$ and the initial Robbins-Monro step size $\gamma_{0}\in\{0.005,0.01,0.05\}$ while keeping the calibration slice fixed at $|\mathcal{C}|=40$.
+For every configuration, we report empirical coverage of the nominal 95% prediction intervals and average interval width (Table 4).
+Table 4: TCP-RM sensitivity on S&P 500 with $|\mathcal{C}|=40$ (target 95%).
+Hyperparameters Performance
+Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
+100 0.005 0.9575 5.3052
+100 0.010 0.9575 5.3066
+100 0.050 0.9581 5.3172
+252 0.005 0.9579 5.1168
+252 0.010 0.9579 5.1191
+252 0.050 0.9586 5.1314
+500 0.005 0.9525 5.4135
+500 0.010 0.9525 5.4319
+500 0.050 0.9567 5.5348
+On S&P 500, w=252w{=}252 yields the narrowest intervals (width $\approx 5.12$), with w=100w{=}100 and w=500w{=}500 producing slightly wider bands and coverage close to or just above the 95% target, except for mild undercoverage in some w=500w{=}500 settings; this motivates our choice w=252w{=}252, $\gamma_{0}{=}0.01$ in the main experiments.
+We also perform sensitivity analyses for TCP-RM on BTC-USD and Gold and TCP (on all three assets), with results reported in Appendix F and Appendix G.
+TCP vs. TCP-RM trace plots are presented in Appendix H.
 7.  Conclusion and Future Work
@@
 Its empirical performance, visual adaptiveness in crisis windows, and robustness to sensible hyperparameters suggest that it is a practical default for risk forecasting and a promising building block for broader sequential decision systems.
-Code Availability
+Code and Data Availability
+The complete implementation of the proposed Temporal Conformal Prediction (TCP) framework is publicly available at https://github.com/agnivibes/temporal-conformal-prediction-tcp/tree/main.
+The repository contains all source code required to reproduce the results reported in this paper, along with the processed financial returns dataset (financial_returns.csv) used in the empirical analysis.
 Appendix A Mathematical Proofs
@@
 ACI 0.5818 2.6643 55
-Appendix C ADDITIONAL TCP VS TCP-RM CRISIS-WINDOW PLOTS
+Appendix C Additional TCP vs TCP-RM Crisis-window Results
 Figure 3 and Figure 4 show the TCP and TCP-RM 95% prediction intervals alongside returns for BTC-USD and Gold during the COVID crisis window (Feb–Apr 2020).
@@
 Appendix F Sensitivity Analysis for TCP-RM
-This section reports the full hyperparameter sensitivity analysis for TCP-RM across the S&P 500, BTC-USD, and Gold.
+This section reports the full hyperparameter sensitivity analysis for TCP-RM across BTC-USD, and Gold.
 For each asset we vary the rolling window size $w\in\{100,252,500\}$ and the initial Robbins-Monro step size $\gamma_{0}\in\{0.005,0.01,0.05\}$ while keeping the calibration slice fixed at $|\mathcal{C}|=40$.
-For every configuration, we report empirical coverage of the nominal 95% prediction intervals and average interval width (Tables 7,8, and 9).
-Table 7: TCP-RM sensitivity on S&P 500 with $|\mathcal{C}|=40$ (target 95%).
-Hyperparameters Performance
-Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
-100 0.005 0.9575 5.3052
-100 0.010 0.9575 5.3066
-100 0.050 0.9581 5.3172
-252 0.005 0.9579 5.1168
-252 0.010 0.9579 5.1191
-252 0.050 0.9586 5.1314
-500 0.005 0.9525 5.4135
-500 0.010 0.9525 5.4319
-500 0.050 0.9567 5.5348
+For every configuration, we report empirical coverage of the nominal 95% prediction intervals and average interval width (Tables 8, and 9).
 Table 8: TCP-RM sensitivity on BTC-USD with $|\mathcal{C}|=40$.
@@
 Summary.
-On S&P 500, w=252w{=}252 yields the narrowest intervals (width $\approx 5.12$), with w=100w{=}100 and w=500w{=}500 producing slightly wider bands and coverage close to or just above the 95% target, except for mild undercoverage in some w=500w{=}500 settings; this motivates our choice w=252w{=}252, $\gamma_{0}{=}0.01$ in the main experiments.
 For BTC-USD, interval widths shrink as $w$ increases, with the tightest bands at w=500w{=}500, while coverage remains stably around 95–96% across all configurations.
@@
 Across all three assets, TCP-RM is therefore stable over the $(w,\gamma_{0})$ grid considered here: the window length $w$ controls the sharpness–stability tradeoff, whereas the specific choice of $\gamma_{0}$ within this range has only a minor effect on coverage and width.
-Appendix G SENSITIVITY ANALYSIS ON TCP (NO RM)
+Appendix G Sensitivity Analysis for TCP (no RM)
 For completeness, in Tables 10, 11 and 12, we also report sensitivity for TCP (split-conformal only, no Robbins–Monro offset). We vary $w\in\{100,252,500\}$ with $|\mathcal{C}|=40$ for each asset.